### v2

# Blown up by an equilateral: Poncelet triangles about the incircle and their degeneracies

Abstract.
We tour several Euclidean properties of Poncelet triangles inscribed in an ellipse and circumscribing the incircle. We also show that a number of degenerate behaviors are triggered by the presence of an equilateral triangle in the family.
1. Introduction
Poncelet’s porism is a 1d family of $n$-gons with vertices on a first conic $\mathcal{E}$ and with sides tangent to a second conic $\mathcal{E}_{c}$ (also called the ‘caustic’). For such a porism to exist, $\mathcal{E}$ and $\mathcal{E}_{c}$ must be positioned in $\mathbb{R}^{2}$ so as to satisfy ‘Cayley’s condition’ [7]. While the porism is projectively invariant – a conic pair is the projective image of two circles where the ‘Poncelet map’ is linearized, – we have found that Poncelet porisms of triangles ($n=3$) are a wellspring of interesting phenomena involving the dynamical geometry of classical objects associated with the triangle (centers, circles, lines/axes, etc.).
An emblematic case, shown in Figure 2 (right) is when when $\mathcal{E},\mathcal{E}_{c}$ are confocal ellipses. In such a porism, – also known as the elliptic billiard – the perimeter is conserved [35], and the loci of all four ‘Greek centers’ of the triangle – incenter, barycenter, circumcenter, orthocenter – sweep ellipses [8, 11, 29]. Significantly, neither foci, perimeter, nor triangle centers are projectively equivariant.
Figure 1. upper left: outer and inner Poncelet conics $\mathcal{E},\mathcal{E}_{c}$ are circles (Chapple’s porism); upper right: outer $\mathcal{E}$ is a circle; bottom left: inner $\mathcal{E}_{c}$ is a circle (in red since focus of present work); bottom right: neither $\mathcal{E}$ nor $\mathcal{E}_{c}$ is a circle. Video
Figure 2. left: the loci of triangle centers $X_{k}$, $k=1,2,3,4$ over Chapple’s porism, studied in [21]. live; right: elliptic loci of said triangle centers over the confocal family, studied in [27]. live
Guided by experiment, we have found that certain types of $\mathcal{E},\mathcal{E}_{c}$ pairs form equivalent classes of Euclidean phenomena. One criterion is whether the pair is concentric and axis-aligned or not. Relevant for the present study, is a subdivision of triangle porisms based on whether either Poncelet conic is a circle or not. Referring to Figure 1, we obtain the following four combinations:
  1. (1)
    Both conics are circles: this is Chapple’s porism [3]. Loci of its triangle centers, e.g., incenter $X_{1}$, barycenter $X_{2}$, circumcenter $X_{3}$, orthocenter $X_{4}$ (the $X_{k}$ notation, used here, follows [18]), have been studied in [13, 14, 21, 26]. See Figure 2 (left).
  2. (2)
    Only $\mathcal{E}$ is a circle. This system, always an affine image of (4), lends itself to a symmetric parametrization which is at the basis of our analysis. Aspects of its geometry and phenomena are analyzed in Appendix C.
  3. (3)
    Only $\mathcal{E}_{c}$ is a circle: here we will tour its manifold Euclidean properties.
  4. (4)
    Neither is circular: Euclidean phenomena are sparser, some are still discernible, e.g., the locus of certain centers are always conics, the locus of barycenter $X_{2}$ and orthocenter $X_{4}$ always homothetic to $\mathcal{E}$ (see below), etc.
For case (3), we start by identifying triangle centers whose loci are conics. In [17] we show that the locus of a triangle center which is a fixed linear combination of the barycenter and the orthocenter is a sufficient condition for that center to sweep a conic. Nevertheless, other centers may still sweep conics for some choices of conic pairs. Still lacking a predictive theory for the latter, we use numerical methods to pinpoint their conic loci, as in [11].
Below we describe many curious dependencies between the shape and foci of certain loci and the geometry of either Poncelet conic. A surprising find has been that the presence of a single equilateral triangle in the family triggers a host of harmonious degeneracies.
1.1. Article structure
We begin with Section 2, describing two algebraic tool-kits used throughout the paper:
  • Section 2.1: we derive the ‘Cayley’ condition for closure of a triangle family interscribed between an ellipse $\mathcal{E}$ and a circle $\mathcal{K}$ (center $C$), and provide an explicit expression for the caustic radius as a function of the ellipse data. We then visualize the quartic $C$-isocurves of the radius of $K$.
  • Section 2.2: we present a parametrization for Poncelet triangles inscribed in the unit circle in $\mathbb{C}$ which keeps the intrinsic symmetry of the vertices, using the work on ‘Blaschke products’ from [4]. We then present the generalization to Poncelet families of convex triangles interscribed between any two ellipses as described in [17], and then specialize it to the case where the inner ellipse is a circle, providing explicit formulas.
These are our main results:
  • Section 3: we characterize remarkable properties of loci of certain triangle centers, including the barycenter, circumcenter, orthocenter, Euler center, and the Gergonne and Nagel points, denoted $X_{2}$, $X_{3}$, $X_{4}$, $X_{5}$, $X_{7}$, and $X_{8}$, respectively, after [18]. The behaviors include:
    • Axis-alignment, concentricity, and homothety with respect to $\mathcal{E}$.
    • The ‘railing’ of the foci of certain centers to the axes of $\mathcal{E}$ or to the caustic center $C$.
    • Circularity: for no obvious reason, some centers always trace circles, regardless of $C$. $X_{36}$, the inversive image of the incenter (stationary in our case) with respect to the (moving) circumcircle is always a circle. We derive expressions for its center and radius.
  • Section 4: A series of curious, ‘degenerate’ phenomena occurs if the Poncelet family contains an equilateral triangle. This is tantamount to $C$ lying on a special ellipse $\mathcal{E}_{eq}$ which is the locus of the centroids of all equilaterals inscribed in $\mathcal{E}$, and derived in [34]. The following facts arise:
    1. (1)
      $C$ is a vertex of the locus of $X_{3}$.
    2. (2)
      The locus of $X_{5}$ collapses to a segment.
    3. (3)
      The Feuerbach point $X_{11}$ becomes stationary on the incircle, and the reflection of the incenter about it ($X_{80}$) is a fixed point on $\mathcal{E}$.
    4. (4)
      The aspect ratio of the loci of the circumcenter $X_{3}$ and the Gergonne point $X_{7}$ are invariant over all $C$ on $\mathcal{E}_{eq}$.
    5. (5)
      The locus of $X_{36}$ becomes a straight line (infinite radius circle).
    6. (6)
      The locus of $X_{59}$, the ‘isogonal conjugate’ of $X_{11}$ (a type of inversive transformation [37]), normally a high-degree curve, collapses to an ellipse touching $\mathcal{E}$ at a special point.
  • Section 5: the envelopes of both the circumcircle and of the (incircle,circumcircle) radical axis are studied. The former is shown to be the union of two circles centered on the foci of the $X_{3}$ locus; the latter is shown to be always a conic, which degenerates to a parabola if $C$ is on the boundary $\mathcal{E}_{eq}$. We conjecture that either envelope has a conic component if and only if the Poncelet family circumscribes a circle.
  • Section 6: by comparing several Poncelet triangle families, we conjecture that the loci of the Gergonne point $X_{7}$ and the Nagel point $X_{8}$ are conics if and only if (i) the caustic is a circle or (ii) the pair is confocal.
  • Section 7: we describe four special triangle porisms when $\mathcal{E}_{c}$ is a circle, that keep key triangle centers stationary: $X_{1}$ (on a focus of $\mathcal{E}$), $X_{2}$, $X_{4}$, and $X_{7}$. We also identify their conservations, some of which generalize to $n>3$ Poncelet porisms.
The following appendices are provided:
  • Appendix A: a table of ‘affine triples’ used in various calculations involving loci, reproduced from [17, Table 1].
  • Appendix B: the loci of several dozen triangle centers are categorized according to one of six remarkable ‘behaviors’: $\mathcal{E}$-homothety, axis-alignment, focus on $C$, major (resp. minor) axis through $C$, and circularity.
  • Appendix C: a geometric analysis of circle-inscribed Poncelet triangles, or case (2) above; we show that phenomena here are closely related to those in (3).
  • Appendix D: computational expressions for the semiaxes of the locus of the Gergonne point $X_{7}$.
  • Appendix E: a table with most symbols used in the article.
An experimental discovery approach “offers the possibility in some instances of formalizing the inductive leaps that the mathematical mind takes when confronted with what seems to be, logically speaking, incomplete evidence” [6, p.210]. This has certainly been the case for us as most phenomena have been discovered via simulation. In some figure captions we include links to videos as well as interactive animations with a custom-built tool [5].
2. Preliminaries
2.1. Cayley closure condition
The Cayley condition specifies whether two ellipses $\{\mathcal{E},\mathcal{E}_{c}\}$ admit a Poncelet family of $n$-gons [7]. Referring to Figure 1 (bottom right), consider a Poncelet triangle family ($n=3$) inscribed in an ellipse centered at the origin with semiaxis lengths $a$ and $b$. Let $(x_{c},y_{c})$ be the center and $a_{c}$ and $b_{c}$ be the semi-axes lengths of a ‘caustic’ ellipse $\mathcal{E}_{c}$ nested within $\mathcal{E}$, and let $\theta_{c}$ be the angle of its major axis with respect to the ‘horizontal’ major axis of $\mathcal{E}$.
In the calculations below, $c_{c}^{2}=a_{c}^{2}-b_{c}^{2}$. Correcting our own expression in [16, formula (3)]:
Proposition 1.
The Cayley condition for the pair $(\mathcal{E},\mathcal{E}_{c})$ to admit an $n=3$ family is given by:
(1) $(a^{4}b_{c}^{4}+a_{c}^{4}c^{4}+b^{4}b_{c}^{4}-2a^{2}b^{2}b_{c}^{4}-2c^{4}a_{c}^{2}b_{c}^{2})c\theta^{4}-8a^{2}b^{2}c_{c}^{2}\,x_{c}y_{c}s\theta c\theta$
$+$ $\left[2c_{c}^{2}(a^{2}+b^{2})(a^{2}y_{c}^{2}-b^{2}x_{c}^{2})+2c_{c}^{2}b^{2}a^{4}+2(b^{2}b_{c}^{4}-a_{c}^{4}c^{2}-b^{4}c_{c}^{2})a^{2}+2b_{c}^{2}(a_{c}^{2}c^{4}-b^{4}b_{c}^{2})\right]c\theta^{2}$
$+$ $a^{4}y_{c}^{4}+b^{4}x_{c}^{4}+2b^{2}(a^{2}a_{c}^{2}-a^{2}b^{2}-b^{2}b_{c}^{2})x_{c}^{2}+2a^{2}b^{2}x_{c}^{2}y_{c}^{2}-2a^{2}(a^{2}a_{c}^{2}+a^{2}b^{2}-b^{2}b_{c}^{2})y_{c}^{2}$
$+$ $(aa_{c}-ab-bb_{c})(aa_{c}+ab-bb_{c})(aa_{c}-ab+bb_{c})(aa_{c}+ab+bb_{c})=0$
where $c\theta=\cos\theta,\;s\theta=\sin\theta$.
Henceforth, let $\delta=\sqrt{(b^{4}+c^{2}y_{c}^{2})(a^{4}-c^{2}x_{c}^{2})}$. Referring to Figure 1 (bottom left), we specialize this to a circular caustic:
Proposition 2.
The radius $r$ for an $N=3$ circular caustic with center at $(x_{c},y_{c})$ is given by:
$r=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}$
Proof.
When the caustic is the circle $(x_{c},y_{c},r)$, the condition in Equation 1 reduces to a biquadratic on $r$:
$c^{4}r^{4}+2(b^{2}c^{2}x_{c}^{2}-a^{2}c^{2}y_{c}^{2}-a^{2}b^{2}(a^{2}+b^{2}))r^{2}+(a^{2}b^{2}-a^{2}y_{c}^{2}-b^{2}x_{c}^{2})^{2}=0$
which has four real roots, or four complex roots. The smaller positive root of the above yields a caustic interior to $\mathcal{E}$. The other positive root $r^{*}>r$, given by:
$r^{*}=\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}+a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}$
yields a circle which completely encloses or intersects $\mathcal{E}$, potentially leading to a complex porism. ∎
As shown in Figure 3, the expression in Proposition 2 prescribes for the interior of $\mathcal{E}$ a foliation of quartic $C$-isocurves of caustic radius.
Figure 3. The interior of $\mathcal{E}$ is foliated by a family of quartic $C$-isocurves of caustic radius, $C=X_{1}$. A particular caustic is shown on the $r=0.35$ isocurve along with one Poncelet triangle.
2.2. Symmetric parametrization
Identifying $\mathbb{R}^{2}$ with $\mathbb{C}$, we derived in [16, Def. 3] the following parameterization of Poncelet triangle families inscribed in $\mathbb{T}$, the unit circle centered at the origin, based on the work in [4] on Blaschke products:
Theorem 1.
For any Poncelet family of triangles inscribed in the unit circle $\mathbb{T}$ and circumscribed about an ellipse with foci $f,g\in\mathbb{D}$ (the unit disk), we can parametrize the elementary symmetric polynomials of its vertices $z_{1},z_{2},z_{3}\in\mathbb{T}$ as
$z_{1}+z_{2}+z_{3}=$ $f+g+\lambda\overline{f}\overline{g}$
$z_{1}z_{2}+z_{2}z_{3}+z_{3}z_{1}=$ $fg+\lambda(\overline{f}+\overline{g})$
$z_{1}z_{2}z_{3}=$ $\lambda$
where $\lambda$ is a parameter that moves along $\mathbb{T}$.
In order to generalize this parameterization for any Poncelet family of triangles interscribed between two ellipses, we simply apply an affine transformation that takes the unit circle centered at the origin to the desired outer ellipse. Without loss of generality, we can assume that the outer ellipse is also centered at the origin and its major axis is aligned with the real axis of $\mathbb{C}$.
As before, let $\mathcal{E}$ and $\mathcal{E}_{c}$ be the outer and inner ellipses of the Poncelet family of triangles and let $a$ and $b$ be the lengths of the semi-major and semi-minor axes of $\mathcal{E}$, respectively, where $\mathcal{E}$ is axis-aligned and centered at the origin. We can then use the transformation $T:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ defined by $T(x,y)=(ax,by)$ to transform the unit circle into $\mathcal{E}$. With a slight abuse of notation, this transformation can also be represented in the complex plane, $T:\mathbb{C}\rightarrow\mathbb{C}$, as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\overline{z}$. Moreover, we have $T^{-1}:\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}$ acting as $T^{-1}(x,y)=(x/a,y/b)$ or alternatively, $T^{-1}:\mathbb{C}\rightarrow\mathbb{C}$ acting on $\mathbb{C}$ as $T^{-1}(z)=\frac{(1/a+1/b)}{2}z+\frac{(1/a-1/b)}{2}\overline{z}$.
Label the vertices of our general family of Poncelet triangles as $z_{1},z_{2},z_{3}\in\mathcal{E}$. Defining $\mathcal{E}_{pre}:=T^{-1}(\mathcal{E}_{c})$ and $z_{i}^{\prime}:=T^{-1}(z_{i})\in\mathbb{T}$ for $i\in\{1,2,3\}$, we find that the triangles $\{z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}\}$ form a Poncelet family interscribed between the unit circle $\mathbb{T}$ and the ellipse $\mathcal{E}_{pre}$. We can then apply the previous parameterization of the elementary symmetric polynomials in $z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}$. This allows us to parameterize any symmetric rational function of the vertices $z_{1},z_{2},z_{3}$ of the general Poncelet triangle family as follows:
Let $P:\mathbb{C}^{3}\rightarrow\mathbb{C}$ be a symmetric rational function. Our objective is to parameterize the point $P(z_{1},z_{2},z_{3})$. Since $P$ is symmetric on its $3$ inputs, the function $P(T(\cdot),T(\cdot),T(\cdot)):\mathbb{C}^{3}\rightarrow\mathbb{C}$ must also be symmetric on its $3$ inputs. Restricting $T$ to the unit circle $\mathbb{T}$, we can write $T:\mathbb{T}\rightarrow\mathcal{E}$ as $T(z):=\frac{(a+b)}{2}z+\frac{(a-b)}{2}\frac{1}{z}$, which is itself a rational function. Thus, $P(z_{1},z_{2},z_{3})=P(T(z_{1}^{\prime}),T(z_{2}^{\prime}),T(z_{3}^{\prime}))$ is a symmetric rational function on $z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}$. Write $P(z_{1},z_{2},z_{3})=\frac{P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}{P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})}$ where $P_{1}$ and $P_{2}$ are symmetric polynomials. By the Fundamental Theorem of Symmetric Polynomials, we can express $P_{1}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ and $P_{2}(z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime})$ as polynomials in $\sigma_{1},\sigma_{2},\sigma_{3}$ where $\sigma_{1}:=z_{1}^{\prime}+z_{2}^{\prime}+z_{3}^{\prime}$, $\sigma_{2}:=z_{1}^{\prime}z_{2}^{\prime}+z_{2}^{\prime}z_{3}^{\prime}+z_{3}^{\prime}z_{1}^{\prime}$, and $\sigma_{3}:=z_{1}^{\prime}z_{2}^{\prime}z_{3}^{\prime}$. Substituting the symmetric parameterization from Theorem 1, we arrive at a parameterization for $P(z_{1},z_{2},z_{3})$ with $\lambda\in\mathbb{T}$ as the parameter, as desired.
With this method, we have proved many results in [16] and [17] such as:
Theorem 2.
In any Poncelet family of triangles, the loci of the barycenter, circumcenter, orthocenter, nine-point center, and the de Longchamps point are always ellipses. More generally, any fixed linear combination of the barycenter $X_{2}$ and the circumcenter $X_{3}$ traces out an ellipse as its locus.
This method is great for proving results such as this, but not so great for finding explicit formulas for the loci. This is due to the fact that everything depends on the foci $f,g$ of the ellipse $\mathcal{E}_{pre}$ which is generally neither of the ellipses $\mathcal{E},\mathcal{E}_{c}$ that define our Poncelet family. However, particularizing the case where the inner ellipse is a circle will allow us to compute such formulas explicitly from the original data of $\mathcal{E}$ and the circle $\mathcal{E}_{c}$.
From now on, assume that $\mathcal{E}_{c}$ is a circle with center $C=x_{c}+iy_{c}$ where $x_{c},y_{c}\in\mathbb{R}$ and radius $r$ given as in Proposition 2.
Since $T^{-1}$ stretches the real and complex axis by factors of $1/a$ and $1/b$, respectively, $\mathcal{E}_{pre}:=T^{-1}(\mathcal{E}_{c})$ is an axis-aligned ellipse with semi-major axis and semi-minor axis equal to $r/b$ and $r/a$, respectively. In fact, it is a $90^{\circ}$-rotated copy of the outer ellipse $\mathcal{E}$ shrunken by a factor of $r/(ab)$. Therefore its semi-focal length is given by $r\frac{c}{ab}$. Moreover, since ellipse centers are preserved by linear transformations, the center of $\mathcal{E}_{pre}$ is $T^{-1}(C)=x_{c}/a+iy_{c}/b$. Since the major axis of $\mathcal{E}_{pre}$ is parallel to the imaginary axis of $\mathbb{C}$, its foci $f$ and $g$ will be given by $x_{c}/a+iy_{c}/b\pm r\frac{c}{ab}$. In order to use the parameterization from Theorem 1 we explicitly calculate their sum and product as
$f+g=$ $\frac{2x_{c}}{a}+\frac{2y_{c}}{b}$
$fg=$ $\frac{a^{2}+b^{2}}{c^{2}}+\frac{2i}{ab}\left(x_{c}y_{c}+\frac{1}{c}\sqrt{(a^{4}-c^{2}x_{c}^{2})(b^{4}-c^{2}y_{c}^{2})}\right)$
We can now calculate explicit formulas for any Poncelet family of triangles where the inner ellipse is a circle. In particular, we can calculate the loci of the barycenter $X_{2}$ and the circumcenter $X_{3}$, and then use the table in Appendix A to explicitly calculate the loci of many triangle centers that can be conveniently written as a linear combination of $X_{1}$, $X_{2}$, and $X_{3}$.
3. Remarkable loci
In this section, let $\mathcal{T}$ (resp. $\mathcal{T}_{g}$) denote a family of Poncelet triangles circumscribing a circle (resp. interscribed between two generic conics $\mathcal{E},\mathcal{E}_{c}$). In both cases, let the center of the outer conic be the origin and $C=[x_{c},y_{c}]$ denote the caustic’s center.
In [17, Thm.1] we showed that if a triangle center $X$ is a fixed linear combination of the barycenter $X_{2}$ and the circumcenter $X_{3}$, its locus over any Poncelet family is a conic. Since over $\mathcal{T}$ the incenter $X_{1}$ is fixed:
Corollary 1.
Any center which is a fixed linear combination of the incenter $X_{1}$ and a another center which sweeps a conic locus will also sweep a conic as its locus.
3.1. Loci of X(k), k=2,4,7
The barycenter $X_{2}$ (resp. orthocenter $X_{4}$) is the meet-point of lines from each vertex to the opposite side’s midpoint (resp. altitude foot). The Gergonne point $X_{7}$ is the meet-point of lines from each vertex to the where the opposite side touches the inscribed circle. For more information, see [37].
For next three propositions, refer to Figure 4.
Figure 4. Close-up of the first quadrant of $\mathcal{E}$, showing a Poncelet triangle (blue) and the circular caustic centered on $C=X_{1}$. Also shown are the loci of the barycenter $X_{2}$ (red), orthocenter center $X_{4}$ (orange), and Gergonne point $X_{7}$ (green). Video
Proposition 3.
Over $\mathcal{T}$, the locus of $X_{2}$ is homothetic to $\mathcal{E}$. Its center $C_{2}$ is collinear with the center of $\mathcal{E}$ and the caustic center $C$. The center $C_{2}$ and semi-axis lengths $a_{2},b_{2}$ are given by:
$C_{2}$ $=\frac{2}{3}\left[x_{c},y_{c}\right]$
$a_{2}^{2}$ $={\frac{-4\,ab\left({a}^{2}+{b}^{2}\right)\delta-4\,{b}^{4}{c}^{2}x_{c}^{2}+4\,{a}^{4}{c}^{2}y_{c}^{2}+{a}^{2}{b}^{2}\left({a}^{4}+6\,{a}^{2}{b}^{2}+{b}^{4}\right)}{9\,{b}^{2}{c}^{4}}}$
$b_{2}^{2}$ $=\frac{b^{2}}{a^{2}}a_{2}^{2}$
Note that $a_{2}/b_{2}=a/b$, as claimed.
We note that this is a special case of [30, Thm.1], which states that for any pair of conics admitting a Poncelet family, the loci of both vertex and area centroids sweep conics. For triangles, these are coincident.
Referring to Figure 5, the following has been proved in [15]:
Theorem 3.
Over $\mathcal{T}_{g}$ the locus of $X_{4}$ is a conic homothetic to a $90^{o}$-rotated copy of $\mathcal{E}$. The locus center $C_{4}$ only depends on $a,b$ of $\mathcal{E}$ and $C$ and is given by:
$C_{4}$ $=\left[\frac{(a^{2}+b^{2})}{a^{2}}x_{c},\frac{(a^{2}+b^{2})}{b^{2}}y_{c}\right]$
Proposition 4.
Over $\mathcal{T}$ the semiaxis lengths $a_{4},b_{4}$ of the $X_{4}$ locus are given by:
$a_{4}^{2}=\frac{z_{4}}{a^{2}b^{4}c^{4}},\,\,\,b_{4}^{2}=\frac{z_{4}}{b^{2}a^{4}c^{4}}$
where:
$z_{4}$ $=(a^{2}+b^{2})(a^{8}y_{c}^{2}+b^{8}x_{c}^{2}-4a^{3}b^{3}\delta)+a^{8}b^{4}+a^{6}b^{4}\left(6b^{2}-z_{4}^{\prime}\right)+a^{4}b^{6}\left(b^{2}-z_{4}^{\prime}\right)$
$z_{4}^{\prime}$ $=x_{c}^{2}+y_{c}^{2}$
Proof.
CAS. Notice that $a_{4}/b_{4}=a/b$. as required. ∎
Figure 5. left: a Poncelet triangle (blue) interscribed between an outer ellipse $\mathcal{E}$ and a generic caustic $\mathcal{E}_{c}$, let $C=$ be its center. Over the family, the locus of $X_{4}$ (orange) is homothetic to a $90^{o}$-rotated copy of $\mathcal{E}$. right: another caustic centered at $C$ which closes Poncelet. Notice that the center of the (new) locus of $X_{4}$ remains at the same location. Video
Let $\mathcal{L}_{7}$ denote the locus of the Gergonne point $X_{7}$.
Proposition 5.
Over $\mathcal{T}$, $\mathcal{L}_{7}$ is an ellipse with major (resp. minor) axis is parallel to $\mathcal{E}$’s minor (resp. major axis). The coordinates of its center $C_{7}=[x_{7},y_{7}]$ are given by:
$x_{7}$ $=\frac{4ax_{c}\left(4a^{5}+a^{3}b^{2}+ab^{4}-a\left(4a^{2}-3b^{2}\right)x_{c}^{2}-ab^{2}y_{c}^{2}-3b\delta\right)}{a^{2}\left(4a^{2}-b^{2}\right)^{2}-\left(4a^{2}-3b^{2}\right)^{2}x_{c}^{2}-a^{2}b^{2}y_{c}^{2}}$
$y_{7}$ $=\frac{4by_{c}\left(4b^{5}+a^{2}b^{3}+a^{4}b-b\left(4b^{2}-3a^{2}\right)y_{c}^{2}-a^{2}bx_{c}^{2}-3a\delta\right)}{b^{2}\left(4b^{2}-a^{2}\right)^{2}-\left(3a^{2}-4b^{2}\right)^{2}y_{c}^{2}-a^{2}b^{2}x_{c}^{2}}$
Furthermore, the ratio of semi-axis lengths $a_{7}/b_{7}$ is given by:
$\left(\frac{a_{7}}{b_{7}}\right)^{2}=\frac{y_{7}\,x_{c}}{x_{7}\,y_{c}}$
Proof.
Symmetric parametrization yields an axis-aligned conic. ∎
Observation 1.
The semi-axis’ lengths $a_{7},b_{7}$ of $\mathcal{L}_{7}$ are rather long expressions, appearing in Appendix D.
3.2. Loci of X(k), k=3,5,8
The circumcenter $X_{3}$ is the center of the circumscribed circle. The Euler center $X_{5}$ is the center of the nine-point circle, which passes through the sides’ midpoints. The Nagel point $X_{8}$ is the meet-point of lines from each vertex to where the opposite touches an excircle. For more information, see [37].
For the next three propositions, refer to Figure 6. Let $\mathcal{L}_{3}$ be the (elliptic) locus of $X_{3}$ over $\mathcal{T}$.
Figure 6. A Poncelet triangle (blue) is shown inscribed in an outer ellipse and circumscribing the incircle centered at $C=X_{1}$. Also shown are the loci of the circumcenter $X_{3}$ (red), Euler center $X_{5}$ (orange), and Nagel point $X_{8}$ (green). Video 1, Video 2
Proposition 6.
The foci $F_{3},F_{3}^{\prime}$ of $\mathcal{L}_{3}$ are collinear with $C$ and are each railed to an axis of $\mathcal{E}$, namely:
$F_{3}=\left[x_{c}(1-(b/a)^{2}),0\right],\,\,\,F_{3}^{\prime}=\left[0,y_{c}(1-(a/b)^{2})\right]$
Furthermore, the locus semi-axis lengths $a_{3},b_{3}$ of $\mathcal{L}_{3}$ are given by:
(2) $a_{3}=\delta_{3}(a/b)-\delta_{3}^{\prime}(b/a),\,\,\,b_{3}=\delta_{3}^{\prime}-\delta_{3}$
where $\delta_{3}=\frac{\sqrt{b^{4}+c^{2}y_{c}^{2}}}{2b}$ and $\delta_{3}^{\prime}=\frac{\sqrt{a^{4}-c^{2}x_{c}^{2}}}{2a}$.
Observation 2.
In the special case when the circular caustic is concentric with $\mathcal{E}$, (i) $r=(ab)/(a+b)$, (ii) the invariant circumradius $R=(a+b)/2$, and (iii) $X_{3}$ sweeps a circle of radius $(a-b)/2$ concentric with $\mathcal{E}$ [9, Section 3]. If the latter’s center is thought of as two coinciding foci, note that each is still on an axis of $\mathcal{E}$, as required by Proposition 6.
Let $\mathcal{L}_{5}$ denote the (elliptic) locus of $X_{5}$ over $\mathcal{T}$.
Proposition 7.
Ae focus $F_{5}$ of $\mathcal{L}_{5}$ is at $C=(x_{c},y_{c})$ and the other $F^{\prime}_{5}$ is given by:
$F_{5}^{\prime}=\left[x_{c}\frac{1+(b/a)^{2}}{2},y_{c}\frac{1+(a/b)^{2}}{2}\right]$
Furthermore, thesemi-axis lengths $a_{5},b_{5}$ of $\mathcal{L}_{5}$ are given by:
$a_{5}=\frac{\left(a^{5}+3a^{3}b^{2}\right)\delta_{3}-\left(b^{5}+3a^{2}b^{3}\right)\delta_{3}^{\prime}}{abc^{2}},\,\,\,b_{5}=\frac{\left(3a^{3}+ab^{2}\right)\delta_{3}-\left(3b^{3}+a^{2}b\right)\delta_{3}^{\prime}}{c^{2}}$
where $\delta_{3},\delta_{3}^{\prime}$ are as defined for the locus of $X_{3}$.
Proposition 8.
Over $\mathcal{T}$, the locus of the Nagel point $X_{8}$ is an ellipse homothetic and concentric with $\mathcal{E}$. Its semi-axis lengths $a_{8},b_{8}$ are given by:
$a_{8}=\frac{\sqrt{\delta_{8}}}{bc^{2}},\,\,\,b_{8}=\frac{\sqrt{\delta_{8}}}{ac^{2}}$
where: $\delta_{8}=(a^{3}b+ab^{3}-2\delta)^{2}+4c^{4}x_{c}^{2}y_{c}^{2}$. Note that $a_{8}/b_{8}=a/b$ as claimed.
Proof.
Symmetric parametrization yields a concentric, axis-aligned conic. ∎
3.3. Circular X(36)
The inversive image of the incenter $X_{1}$ with the respect to the circumcircle is called $X_{36}$ on [18]. Referring to Figure 7, the following is a rather surprising phenomenon:
Proposition 9.
Over $\mathcal{T}$, the locus of $X_{36}$ is a circle. Its center $C_{36}$ is given by:
$C_{36}$ $=[x_{36}/{z_{1}},y_{36}/{z_{1}}]$
$x_{36}$ $=x_{c}\left({z_{2}}-6a^{4}b^{4}+6a^{4}b^{2}y_{c}^{2}-3a^{2}b^{6}-2a^{2}b^{4}x_{c}^{2}+3a^{2}b^{4}y_{c}^{2}-8ab^{3}\,{\delta}+3b^{6}x_{c}^{2}\right)$
$y_{36}$ $=y_{c}\left(3a^{6}\left(y_{c}^{2}-b^{2}\right)-a^{4}\left(6b^{4}+b^{2}\left(2y_{c}^{2}-3x_{c}^{2}\right)\right)-8a^{3}b\,{\delta}+a^{2}b^{4}\left(b^{2}+6x_{c}^{2}-y_{c}^{2}\right)-9b^{6}x_{c}^{2}\right)$
where:
${z_{1}}$ $={z_{2}}-2a^{4}b^{4}-6a^{4}b^{2}y_{c}^{2}+a^{2}b^{6}-6a^{2}b^{4}x_{c}^{2}-a^{2}b^{4}y_{c}^{2}-9b^{6}x_{c}^{2}$
${z_{2}}$ $=a^{6}b^{2}-9a^{6}y_{c}^{2}-a^{4}b^{2}x_{c}^{2}$
Furthermore, the radius $r_{36}$ of the locus $X_{36}$ is given by:
$r_{36}$ $=\frac{2(\delta_{4}z_{36}-\delta_{4}^{\prime}z^{\prime}_{36})}{a^{2}b^{2}c^{4}-b^{2}(a^{2}+3b^{2})^{2}x_{c}^{2}-a^{2}(3a^{2}+b^{2})^{2}y_{c}^{2}}$
$z_{36}$ $=5ba^{4}y_{c}^{2}-b^{3}a^{2}(a^{2}+y_{c}^{2}-x_{c}^{2})+b^{5}(a^{2}+3x_{c}^{2})$
$z^{\prime}_{36}$ $=5ab^{4}x_{c}^{2}-a^{3}b^{2}(b^{2}+x_{c}^{2}-y_{c}^{2})+a^{5}(b^{2}+3y_{c}^{2})$
where $\delta_{4}=\sqrt{a^{4}-c^{2}x_{c}^{2}}$ and $\delta_{4}^{\prime}=\sqrt{b^{4}+c^{2}y_{c}^{2}}$.
Figure 7. The locus of $X_{36}$, the inversion of $C=X_{1}$ (fixed) with respect to the (moving) circumcircle (dashed blue), is a circle (orange). Also shown is the elliptic locus of the circumcenter $X_{3}$ (red), whose foci $F_{3},F_{3}^{\prime}$ lie on the semiaxes of the outer ellipse. Indeed, $F_{3},F_{3}^{\prime},C,C_{36}$ are collinear. Note that from the inversion operation, $X_{3}X_{1}X_{36}$ are dynamically collinear. Indeed, the foci live
Referring to Figure 8:
Proposition 10.
The major axis of $\mathcal{L}_{3}$ passes through the center $C_{36}$ of the $X_{36}$ locus. Furthermore, the circumradius is minimized (resp. maximized) when $X_{3}$ is on a major vertex of its locus closest (resp. farthest) from $C=X_{1}$. At these configurations, the distance $|X_{1}X_{36}|$ maximized (resp. minimized), respectively.
Proof.
Direct from Proposition 6 and Proposition 9. ∎
Figure 8. The major axis of the $X_{3}$ locus passes through the center $C_{36}$ of the circular locus of $X_{36}$. On the left (resp. right), $X_{3}$ is at the major vertex of the locus closest to (resp. farthest from) $C=X_{1}$, in which case the circumradius is minimized (resp. maximized) and $X_{36}$ is farthest from (resp. closest to) $C$. live
Figure 9. With $C=X_{1}=[0,0]$, the locus of $X_{36}$ (red) is a circle concentric with $\mathcal{E}$. In the figure, $a=7/2$ and $b=1$, so $r_{36}=14/5$. Also shown are the loci of $X_{2}$ (blue ellipse), $X_{3}$ (green circle) and $X_{4}$ (purple ellipse). live
Referring to Figure 9:
Corollary 2.
Over $\mathcal{T}$, when $C$ is at the center of $\mathcal{E}$, the locus of $X_{36}$ is a circle concentric with $\mathcal{E}$, $r_{36}$ reduces to:
$r^{*}_{36}=\frac{2ab}{a-b}$
Proof.
Direct from setting $x_{c}=y_{c}=0$ in $r_{36}$ (Proposition 9). ∎
4. Degeneracies
This section described many degeneracies brought about by the ‘presence’ of an equilateral triangle in $\mathcal{T}$. The locus of centroids of a 1d family of equilateral triangles inscribed in an ellipse $\mathcal{E}=(a,b)$ is an ellipse – henceforth called $\mathcal{E}_{eq}$ – concentric and axis-aligned with $\mathcal{E}$, whose semi-axes $a_{eq},b_{eq}$ are given by [32, 34]:
$a_{eq}=\frac{a\,c^{2}}{a^{2}+3b^{2}},\,\,\,b_{eq}=\frac{b\,c^{2}}{3a^{2}+b^{2}}.$
Let $\mathcal{T}^{*}$ denote the family of Poncelet triangles about a circle with center $C=[a_{eq}\cos t,b_{eq}\sin t]$ on $\mathcal{E}_{eq}$.
Corollary 3.
$\mathcal{T}^{*}$ contains an equilateral triangle.
Let $V^{*}=[a\cos(t^{*}),b\sin(t^{*})]$ on $\mathcal{E}$ be a vertex of said equilateral.
Proposition 11.
$t^{*}$ is given by the three solutions of:
$(a^{2}+3b^{2})\left[(3a^{2}+b^{2})\cos^{2}t^{*}-2b^{2}\sin^{2}t-2b^{2}\sin t\sin t^{*}\right]$
$+(3a^{2}+b^{2})\left[-2a^{2}\cos t\cos t^{*}+(a^{2}-3b^{2})\cos^{2}t\right]=0$
Proof.
It can be shown that a triangular orbit will be equilateral when the triangle $\{V^{*},C,[a\cos t,-b\sin t]\}$ is isosceles. ∎
Sample loci of $X_{k}$, $k=2,3,4,5$ over $\mathcal{T}^{*}$ are shown Figure 10 (top). Notice they simultaneously pass through $X_{1}$ when the Poncelet triangle becomes an equilateral. In Figure 10 (bottom) are shown the loci of $k=7,8,11,80$. The first two display the same behavior, while the latter two are stationary.
Figure 10. top: when the incenter is on $\mathcal{E}_{eq}$ (dark green), the loci of $X_{k}$, $k=2,3,4,5$ (red, light green, blue, purple), pass simultaneously through the incenter when the equilateral in the family is traversed, notice that the locus of $X_{5}$ (purple) has collapsed to a segment, live; botom: same situation, showing the loci of $X_{k}$, $k=7,7,11,80$ (red, light green, blue, purple). The latter two are stationary on the incircle and $\mathcal{E}$, respectively, and along the $X_{1}X_{5}$ line, live
4.1. C is a vertex of X(3)
Referring to Figures 11 and 12:
Corollary 4.
Over $\mathcal{T}^{*}$ $C$ is a (major) vertex of $\mathcal{L}_{3}$. If $C$ is interior (resp. exterior) to $\mathcal{E}_{eq}$, it is interior (resp. exterior) to $\mathcal{L}_{3}$.
Proof.
This stems from the fact that the major axis of the $X_{3}$ locus passes through $C$, Proposition 6. ∎
Figure 11. On the left (resp. right), $C=X_{1}$ is outside (resp. inside) $\mathcal{E}_{eq}$ (green). In this case, $C$ is exterior (resp. interior) to the locus of $X_{3}$ (red), and the locus of $X_{36}$ (orange) is disjoint with (resp. contains) $\mathcal{E}$.
Figure 12. On the left (resp. right), $C=X_{1}$ is on (resp. at the center of) $\mathcal{E}_{eq}$ (green). In this case, $C$ is a vertex (at the center) of the locus of $X_{3}$ (red), and the locus of $X_{36}$ (orange) is a line (resp. a circle concentric with $\mathcal{E}$).
4.2. Invariant eccentricity: X(3), X(7)
Proposition 12.
Over all $C$ on $\mathcal{E}_{eq}$ the aspect ratio of $\mathcal{L}_{3}$ is invariant and given by:
$\frac{a_{3}}{b_{3}}=\frac{\left(3a^{2}+b^{2}\right)\left(a^{2}+3b^{2}\right)\left(a^{4}-b^{4}\right)}{c^{2}\left(6a^{5}b+20a^{3}b^{3}+6ab^{5}\right)}$
Proof.
Evaluating Equation 2 over Proposition 17. ∎
Proposition 13.
Over all $C$ on $\mathcal{E}_{eq}$, the aspect ratio $\mathcal{L}_{7}$ is invariant and given by:
$\frac{a_{7}}{b_{7}}={{\frac{2\,{a}^{2}+{b}^{2}}{\sqrt{2\,{a}^{4}+5\,{a}^{2}{b}^{2}+2\,{b}^{4}}}}}$
Proof.
Direct from Proposition 5. ∎
4.3. Segment X(5)
Proposition 14.
Over all $C=[a_{eq}\cos{t},b_{eq}\sin{t}]$ on $\mathcal{E}_{eq}$, $\mathcal{L}_{5}$ is a segment ($b_{5}=0$) of length $l_{5}$ with an endpoint on $C$ and the other at $F^{\prime}_{5,{eq}}$. These are given by:
$l_{5}^{2}$ $=\frac{c^{8}\left(a^{6}+15a^{4}b^{2}+15a^{2}b^{4}+b^{6}-c^{6}\cos(2t)\right)}{16\left(3a^{5}b+10a^{3}b^{3}+3ab^{5}\right)^{2}}$
$F^{\prime}_{5,{eq}}$ $=\left[\frac{\left(a^{4}-b^{4}\right)\cos{t}}{2\left(a^{3}+3ab^{2}\right)},\frac{\left(a^{4}-b^{4}\right)\sin{t}}{2\left(3a^{2}b+b^{3}\right)}\right]$
Proof.
Applying Proposition 7 on Proposition 17. ∎
Corollary 5.
Over $\mathcal{T}^{*}$ the locus of $X_{12}$ (internal center of similitude of the incircle and the Euler circle [18]), is a segment with one endpoint at $C$.
Proof.
This must be the case because the segment locus of the Euler center $X_{5}$ passes through $C$. ∎
4.4. Stationary X(11), X(80)
Referring to Figure 19:
Proposition 15.
Over $\mathcal{T}^{*}$, the Feuerbach point $X_{11}$ (resp. $X_{80}$, the reflection of $X_{1}$ on $X_{11}$) is stationary at the non-$C$ intersection of the segment-locus of $X_{5}$ with the incircle (resp. with $\mathcal{E}$). With $C=[a_{eq}\cos{t},b_{eq}\sin{t}]$ on $\mathcal{E}_{eq}$, these are given by:
$X^{*}_{11}$ $=\frac{a^{2}+b^{2}}{c^{2}}\left[a_{eq}\cos t,-b_{eq}\sin t\right]$
$X^{*}_{80}$ $=\left[a\cos t,-b\sin t\right]$
Note that $X_{k},k=1$, $5$, $11$, $12$, $80$, …, are points on the $X_{1}X_{5}$ line [19].
Proof.
CAS manipulation. ∎
Observation 3.
Over $\mathcal{T}^{*}$, the locus of $X_{106}$ (isog. conj. of $X_{1}X_{2}$ with the line at infinity) is a circle passing through $X_{80}$.
Lemma 1.
Let $C=[a_{eq}\cos{t},b_{eq}\sin{t}]$. Over $\mathcal{T}^{*}$, the intersection $I^{*}=[x^{*},y^{*}]$ of the ray $X_{1}X_{5}$ with $\mathcal{E}$ is given by:
$x^{*}=$ $\frac{\left(a^{2}v\left(a^{4}-3a^{2}b^{2}-2b^{4}\right)\sin^{2}{t}-b^{4}u^{2}\cos^{2}{t}\right)a\cos{t}}{b^{4}u^{2}\cos^{2}{t}+a^{4}v^{2}\sin^{2}{t}}$
$y^{*}=$ $\frac{\left(b^{2}u\left(2a^{4}+3a^{2}b^{2}-b^{4}\right)\cos^{2}{t}+a^{4}v^{2}\sin^{2}{t}\right)b\sin{t}}{b^{4}u^{2}\cos^{2}{t}+a^{4}v^{2}\sin^{2}{t}}$
where $u=(3\,a^{2}+b^{2})$ and $v=(3\,b^{2}+a^{2})$.
Referring to Figure 13:
Proposition 16.
In $\mathcal{T}^{*}$, the Poncelet triangle $I^{*}B_{1}B_{2}$ is an isosceles, unique in the family (other than the equilateral), with base $B_{1}B_{2}$, the chord of $E$ tangent to the incircle at the stationary $X_{11}$.
Proof.
The triangle is an isosceles since $I^{*},X_{1},X_{5}$ are collinear. The base must be tangent to the incircle (via Poncelet) at $X_{11}$, since the latter is collinear with $X_{1}$ and $X_{5}$ [18]. We omit the expressions for $B_{1}$ and $B_{2}$ since they are rather long. ∎
Figure 13. Construction for Proposition 16: $C=X_{1}$ on $\mathcal{E}_{eq}$, the Poncelet triangle has a vertex on the intersection $I^{*}$ of the ray $X_{1}X_{5}$ with $\mathcal{E}$. The said triangle is an isosceles, and its base $B_{1}B_{2}$ is the chord of $\mathcal{E}$ touching the incircle at $X_{11}$. live
4.5. Line X(36)
In Proposition 9 we saw that over $\mathcal{T}$, the locus $X_{36}$ was a circle. Nevertheless, and referring to Figure 12 (left):
Proposition 17.
Over $\mathcal{T}^{*}$, with $C=[x_{c},y_{c}]$ on $\mathcal{E}_{eq}$, the locus of $X_{36}$ degenerates to the line:
$b^{2}x_{c}x+a^{2}y_{c}y=\frac{b^{2}\left(a^{4}+2a^{2}b^{2}+5b^{4}\right)x_{c}^{2}+a^{2}\left(5a^{4}+2a^{2}b^{2}+b^{4}\right)y_{c}^{2}}{c^{4}}$
Proof.
Direct from Proposition 9, taking the limit when $(x_{c},y_{c})$ is on $\mathcal{E}_{eq}$ and simplifying. ∎
Referring to Figure 11:
Observation 4.
Over $\mathcal{T}^{*}$, the locus of $X_{36}$ is disjoint with (resp. contains) $\mathcal{E}$ if $C$ is exterior (resp. interior) to $\mathcal{E}_{eq}$.
Referring to Figure 14:
Proposition 18.
If $C=[x_{c},y_{c}]$ is on $\mathcal{E}_{eq}$, the major axis of $\mathcal{L}_{3}$ is perpendicular to the line locus of $X_{36}$. These meet at:
$X^{\perp}_{36}=c^{-4}\left[x_{c}(a^{4}+2a^{2}b^{2}+5b^{4}),y_{c}(b^{4}+2b^{2}a^{2}+5a^{4})\right]$
Proof.
Direct from Proposition 6 which the $X_{3}$-locus major axis, and Proposition 17. ∎
Figure 14. When $X_{3}$ is on the non-$X_{1}$ major vertex of its locus (red), $X_{3}X_{1}$ is perpendicular to the line-locus of $X_{36}$. At this moment, $|X_{1}X_{36}|$ is minimal. live
Observation 5.
Over $\mathcal{T}^{*}$, when $X_{3}$ passes through $C=X_{1}$, i.e., the major vertex of its locus, $X_{36}$ is on the line at infinity.
Referring to Figure 15, we specialize Proposition 10 to its degenerate form:
Proposition 19.
Over $\mathcal{T}^{*}$, when $X_{3}$ is on the non-$X_{1}$ major vertex of its locus, the distance $|X_{1}X_{36}|$ is minimal and given by:
$|X_{1}X_{36}|_{min}=\frac{4\left(a^{2}+b^{2}\right)\sqrt{a^{4}y_{c}^{2}+b^{4}x_{c}^{2}}}{c^{4}}$
Furthermore, the circumradius $R$ is minimized (resp. maximized) when $X_{3}=X_{1}$ (resp. $X_{3}$ is on the non-$X_{1}$ vertex of its locus).
$R_{min}$ $=2\left(\frac{b\sqrt{a^{4}-c^{2}x_{c}^{2}}-a\sqrt{b^{4}+c^{2}y_{c}^{2}}}{c^{2}}\right)$
$R_{max}$ $=\frac{\left(a^{2}+b^{2}\right)\sqrt{a^{4}+6a^{2}b^{2}+b^{4}}\,\sqrt{a^{4}y_{c}^{2}+b^{4}x_{c}^{2}}}{2a^{2}b^{2}c^{2}}$
Figure 15. left: When $X_{3}$ is on the non-$X_{1}$ vertex of its locus, the circumradius $R$ is maximized and $|X_{1}X_{36}|$ is minimal; right: if $X_{3}$ on $X_{1}$, the triangle is equilateral, $R$ is minimized while $|X_{1}X_{36}|$ is ‘maximal’, since $X_{36}$ is on the line at infinity. live
Referring to Figure 16:
Observation 6.
Over all $C$ on $\mathcal{E}_{eq}$, the envelope of the line locus of $X_{36}$ is an oval of degree 6.
Figure 16. Over all $C=X_{1}$ on $\mathcal{E}_{eq}$ (dashed green), the envelope of the line locus of $X_{36}$ is a non-conic of degree 6 (dashed orange).
4.6. Elliptic X(59)
Referring to Figure 17, in several Poncelet families studied, the locus of $X_{59}$, the isogonal conjugate of $X_{11}$, is either an oval or a self-intersected curve, e.g., see [11, Fig.2] and [26, Fig.9].
Nevertheless, it collapses to a conic in the following two cases. Firstly:
Proposition 20.
In Chapple’s porism, the locus of $X_{59}$ is an ellipse with major axis on $X_{1}X_{3}$, with semi-axis lengths given by:
$a_{59,chapple}=R,\,\,\,b_{59,chapple}=\frac{R\sqrt{{R}^{2}-{d}^{2}}}{\sqrt{9\,{R}^{2}-{d}^{2}}}$
Figure 17. The locus of $X_{59}$ (isogonal conjugate of $X_{1}$ in (i) the confocal pair (top left), (ii) a generic incircle porism (top right), (iii) Chapple’s porism, where it is an ellipse (top bottom), and (iv) an incircle porism with $X_{1}$ on $\mathcal{E}_{eq}$ (dark green), notice $X_{59}$ is a conic, tilted with respect to $\mathcal{E}$, which it touches at a single point.
The isogonal conjugate $g(P)$ of a point $P$ with respect to a triangle $T$ is the point of concurrence of the cevians of $P$ reflected upon the angle bisectors [37]. In [33] it is shown that the locus of $g(P$) over circle-inscribed Poncelet triangles is a circle. As above, let $\mathcal{T}_{g}$ be a family of Poncelet triangles interscribed between two generic conics $\mathcal{E},\mathcal{E}_{c}$. Referring to Figure 18, the following is proved in [10]:
Lemma 2.
Over $\mathcal{T}_{g}$, the locus of $g(P)$ is a conic. If $P$ is on $\mathcal{E}_{c}$ the locus is an ellipse interior to $\mathcal{E}$ and touching it at a point. If $P$ is on $\mathcal{E}$ the locus is the union of a straight line and the line at infinity (degenerate hyperbola).
An alternative proof to Proposition 17 is to based on the fact that $X_{36}$ and $X_{80}$ are isogonal conjugates [18]:
Corollary 6.
Over $\mathcal{T}^{*}$, the locus of $X_{36}$ is a line since $X_{80}$ is stationary on $\mathcal{E}$, Proposition 15.
Figure 18. Results from [10]: left: over generic Poncelet triangles, the locus of the isogonal conjugate $g(P)$ of a fixed point $P$ is a conic (purple); right: if $P$ is on the caustic, the locus of $g(P)$ is an ellipse that touches $\mathcal{E}$ at one point. Video
Secondly, referring to Figure 17 (bottom right) and Figure 19:
Corollary 7.
Over $\mathcal{T}^{*}$, the locus of $X_{59}$ is an ellipse interior to $\mathcal{E}$ and touching it at a point.
Proof.
$X_{59}$ is the isogonal conjugate of $X_{11}$ [18], which is stationary on the incircle over $\mathcal{T}^{*}$. ∎
Furthermore, let $I^{*}$ denote the touch-point between the elliptic locus of $X_{59}$ (over $\mathcal{T}^{*}$) and $\mathcal{E}$.
Observation 7.
When the Poncelet triangle is an isosceles, i.e., one of its vertices is $I^{*}$, $X_{59}$ is at $I^{*}$. When the Poncelet triangle is an equilateral, $X_{59}$ ‘blows up’ at a major vertex of its locus.
Figure 19. When the incenter is on $\mathcal{E}_{eq}$ (dark green), the locus of $X_{5}$ is a line, and $X_{11}$ and $X_{80}$ are stationary on the incircle and $\mathcal{E}$, respectively. The locus of $X_{59}$ is an ellipse (blue), internally-tangent to $\mathcal{E}$ at $I^{*}$ (see Lemma 1) In the picture, $X_{5}$ is on $X_{1}$, i.e., the Poncelet triangle is an instantaneous equilateral. When this happens, $X_{59}$ is at a vertex of its locus. live
5. Envelopes
As before, let $\mathcal{T}$ be a family of Poncelet triangles inscribed in an ellipse $\mathcal{E}$ and circumscribing a circle $\mathcal{K}$ with center $C=X_{1}$.
5.1. Circumcircle
Referring to Figure 20:
Proposition 21.
Over $\mathcal{T}$, the envelope of the circumcircle are two nested circles, each tangent to $\mathcal{E}$ at two points and centered at the foci $F_{3}$ and $F_{3}^{\prime}$ of $X_{3}$ locus (Proposition 6).
Proof.
Let $C=(x_{c},y_{c})$ be the center of caustic and $r$ the radius of the caustic given by Proposition 2; the incircle $\mathcal{K}$ is then given by $(x-x_{c})^{2}+(y-y_{c})^{2}=r^{2}$. Let $\lambda=\cos{u}+i\sin{u}$ in the symmetric parametrization of Section 2.2, obtain the circumcircle $\mathcal{K}^{\prime}$:
$\mathcal{K}^{\prime}(x,y)=x^{2}+y^{2}-\left(\frac{\left(a^{3}b-\delta\right)\cos{u}}{b\,a^{2}}+\frac{c^{2}y_{c}x_{c}\sin{u}}{b\,a^{2}}+\frac{c^{2}x_{c}}{a^{2}}\right)x$
$+$ $\left(-\frac{c^{2}x_{c}y_{c}\cos{u}}{a\,b^{2}}+\frac{c^{2}y_{c}}{b^{2}}+\frac{\left(a\,b^{3}-\delta\right)\sin{u}}{a\,b^{2}}\right)y+\frac{c^{2}x_{c}\cos{u}}{a}+\frac{c^{2}y_{c}\sin{u}}{b}-\frac{\delta}{ba}=0$
The envelope of this family of circles is obtained by setting:
$\mathcal{K}^{\prime}(x,y)=0,\,(\partial{\mathcal{K}^{\prime}}/\partial u)(x,y)=0$
With the above equations this yields the union of two circles $\mathcal{K}_{1}=(O_{1},r_{1})$ and $\mathcal{K}_{2}=(O_{2},r_{2})$ where:
$O_{1}=$ $[0,-y_{c}c^{2}/b^{2}]=[0,y_{c}(1-(a/b)^{2})]=F_{3}^{\prime}$
$r_{1}=$ $(a/b^{2})\sqrt{b^{4}+c^{2}y_{c}^{2}}$
$O_{2}=$ $[x_{c}c^{2}/a^{2},0]=[x_{c}(1-(b/a)^{2}),0]=F_{3}$
$r_{2}=$ $(b/a^{2})\sqrt{a^{4}-c^{2}x_{c}^{2}}$
$\mathcal{K}_{1}$ touches $\mathcal{E}$ at [±$\left[\pm(a/b)\sqrt{b^{2}-y_{c}^{2}},y_{c}\right]$; $\mathcal{K}_{2}$ touches $\mathcal{E}$ at $\left[x_{c},\pm(b/a)\sqrt{a^{2}-x_{c}^{2}}\right]$. ∎
Figure 20. The envelope of the circumcircle consists of two nested circles twice tangent to $\mathcal{E}$ and centered on the foci of the $X_{3}$ locus (red). Left, middle, right show three positions for $C=X_{1}$: interior, on the boundary, and exterior to $\mathcal{E}_{eq}$ (green), respectively. Video
Given two fixed nested (resp. unnested) circles, the locus of the center of circles simultaneously tangent to them is a conic with foci on their centers and with major axis length equal to the difference (resp. sum) of their radii [22, thm.14, p.57]. In our case, the circumcircles of a Poncelet family circumscribing a circle are tangent to the two circular boundaries of their envelope. In Equation 2, an expression is derived for $a_{3}$, the semi-axis length of the locus of $X_{3}$. It follows that:
Corollary 8.
$|r_{1}-r_{2}|=2a_{3}$.
For the general case, experimental evidence suggests:
Conjecture 1.
Over $\mathcal{T}_{g}$, either component of the envelope of the circumcircle is a conic if and only if $\mathcal{E}_{c}$ is a circle, in which case both components are circles (Proposition 21).
Note that if the outer conic is a circle, said envelope is not defined.
5.2. Radical axis
Let $\mathcal{R}$ denote the radical axis of incircle and circumcircle of a triangle, i.e., the line whose points have equal circle power with respect to the two said circles [37]. $X_{1}X_{3}$ must perpendicular to $\mathcal{R}$ since it is the line that passes through both circle centers. According to [18], $X_{3660}$ is the ‘trace’ of $\mathcal{R}$, i.e., the perpendicular intersection of $X_{1}X_{3}$ with the radical axis.
Referring to Figure 21:
Proposition 22.
Over $\mathcal{T}$, the envelope of $\mathcal{R}$ is a conic whose major axis coincides with that of the $X_{3}$ locus (see Proposition 6). It will be an ellipse (resp. hyperbola) if $C$ is interior (resp. exterior) to $\mathcal{E}_{eq}$.
Proof.
Let $C=(x_{c},y_{c})$ be the center of caustic and $r$ the radius of the caustic given by Proposition 2; the incircle $\mathcal{K}$ is then given by $(x-x_{c})^{2}+(y-y_{c})^{2}=r^{2}$. Let $\lambda=\cos{u}+i\sin{u}$ in the symmetric parametrization of Section 2.2, obtain the circumcircle $\mathcal{K}^{\prime}$:
$\mathcal{K}^{\prime}(x,y)=x^{2}+y^{2}-\left(\frac{\left(a^{3}b-\delta\right)}{b\,a^{2}}\cos u+\frac{c^{2}y_{c}x_{c}}{b\,a^{2}}\sin{u}+\frac{c^{2}x_{c}}{a^{2}}\right)x$
$+$ $\left(-\frac{c^{2}x_{c}y_{c}}{a\,b^{2}}\cos{u}+\frac{\left(a\,b^{3}-\delta\right)}{a\,b^{2}}\sin{u}+\frac{c^{2}y_{c}}{b^{2}}\right)y+\frac{c^{2}x_{c}}{a}\cos{u}+\frac{c^{2}y_{c}}{b}\sin{u}-\frac{\delta}{ba}=0$
The radical axis of $(K,\mathcal{K}^{\prime})$, is given by $l(u)x+m(u)y+n(u)=0$, with:
$l(u)$ $=c^{4}\left(-bx_{c}y_{c}\,c^{2}\sin{u}+b\left(\delta-a^{3}b\right)\cos{u}+b^{2}\left(a^{2}+b^{2}\right)x_{c}\right)$
$m(u)$ $=c^{4}\left(a\left(a\,b^{3}-\delta\right)\sin{u}-ax_{c}y_{c}\,c^{2}\cos{u}+a^{2}\left(a^{2}+b^{2}\right)y_{c}\right)$
$n(u)$ $=a^{2}by_{c}\,c^{6}\sin{u}+a\,b^{2}x_{c}\,c^{6}\cos{u}+a^{2}b^{2}c^{2}\left(-a^{2}x_{c}^{2}+b^{2}y_{c}^{2}\right)+a^{4}b^{4}\left(a^{2}+b^{2}\right)$
$-ab\left(a^{4}+b^{4}\right)\delta.$
The envelope of the family of lines above is given by:
$E(u)=\left[\frac{n^{\prime}(u)m\left(u\right)-m^{\prime}(u)n\left(u\right)}{m^{\prime}(u)l\left(u\right)-l^{\prime}(u)m\left(u\right)},-\frac{n^{\prime}(u)l\left(u\right)-l^{\prime}(u)n\left(u\right)}{m^{\prime}(u)l\left(u\right)-l^{\prime}(u)m\left(u\right)}\right]$
It follows from the calculations that the envelope is parametrized by:
$E(u)=\left[\frac{k_{1}\sin u+k_{2}\cos u+k_{3}}{d_{1}\sin u+d_{2}\cos u+d_{3}},\frac{m_{1}\sin u+m_{2}\cos u+m_{3}}{d_{1}\sin u+d_{2}\cos u+d_{3}}\right]$
Here all constants are long expressions involving the variables $(a,b,c,x_{c},y_{c}).$ Eliminating the variables we obtain that the envelope is given implicitly by a quadratic equation and so it is a conic. Algebraic manipulation yields its major axis to be as the one in Proposition 6. ∎
Referring to Figure 22:
Corollary 9.
Over $\mathcal{T}$, the envelope of $\mathcal{R}$ is a parabola if $C$ is on the boundary of $\mathcal{E}_{eq}$ whose axis of symmetry is the major axis of the $X_{3}$ locus, i.e., its directrix is parallel to the line-locus of $X_{36}$ (Proposition 17).
Proof.
When the family is at the equilateral position, $X_{1}$ and $X_{3}$ coincide, and $\mathcal{R}$ is the line at infinity. ∎
Figure 21. Left (resp. right): envelope (orange) of $\mathcal{R}$ (dashed magenta) is an ellipse (resp. hyperbola), when $C=X_{1}$ is interior (resp. exterior) to $\mathcal{E}_{eq}$ (green ellipse). The major axis of the locus of $X_{3}$ (red) coincides with that of the envelopes.
Figure 22. When $C=X_{1}$ is on the boundary of $\mathcal{E}_{eq}$ (green ellipse), the envelope of $\mathcal{R}$ (dashed magenta) is parabolic (orange, focus $F$, vertex $V$); the axis of symmetry (dashed orange) is the major axis of the locus of $X_{3}$ (red). Also shown is the directrix of the parabola (dashed orange), with $V^{\prime}$ its intersection with the axis of symmetry.
Based on experimental evidence:
Conjecture 2.
Over $\mathcal{T}$, the envelope of $\mathcal{R}$ is a conic if and only if the Poncelet caustic is a circle.
The Feuerbach point $X_{11}$ is where incircle and nine-point circle touch. Therefore, the radical axis of said circles is tangent to the incircle at $X_{11}$. Let said axis be called $L(101)$ after [37]. Therefore:
Corollary 10.
Over $\mathcal{T}$, the envelope of $L(101)$ is the incircle itself. If $C$ is on $\mathcal{E}_{eq}$, $X_{11}$ is stationary (Proposition 15), therefore, in such a case said envelope is undefined.
The orthic axis, called $\mathcal{L}_{3}$ in [37] (resp. anti-orthic axis, $\mathcal{L}_{1}$) is the radical axis of the circumcircle and the nine-point (resp. Bevan) circle [37, Radical line], the latter is centered on $X_{5}$ (resp. $X_{40}$) and has radius $R/2$ (resp. $2R$). Experimentation shows:
Observation 8.
Over $\mathcal{T}$, the envelope of both $\mathcal{L}_{3}$ and $\mathcal{L}_{1}$ are conics: an (ellipse, parabola, hyperbola) if $C$ is (interior, on the boundary, exterior) to $E_{eq}$ respectively. Nevertheless, the envelope axes and those of the $X_{3}$ locus are not axis-aligned.
6. Ellipticity of X(k), k=1, 7, 8
In [11, 29] it is proved that the locus of the incenter $X_{1}$ is a conic if the Poncelet ellipse pair is confocal (elliptic billiard). In [16, Conj.3] it is conjectured that $X_{1}$ sweeps a conic if and only if the pair is confocal, and a proof is still outstanding.
Experiments with ten different Poncelet families (see six of them in Figures 23 and 24) are summarized in Table 1. In turn, these suggest:
Conjecture 3.
The locus of $X_{7}$ and/or $X_{8}$ are conics if and only if (i) the Poncelet conic pair is confocal or (ii) the caustic is a circle (in which case $X_{1}$ is stationary).
Family Outer Inner $X_{1}$ $X_{7}$ $X_{8}$ bit.ly/<code>
Confocal E E E E E 46KYk4D
Chapple C C P C C 4cqyGmN
Incircle E C P E E 3M50bHG
Circ-caustic E C P E E 4fK8Exy
Homothetic E E - - - 3M4sG8G
Dual E E - - - 3WHkQqg
Conf. Excentrals E E - - - 3WGARgk
Inellipse C E - - - 4fKZPUv
Brocard C E - - - 4fKQmMH
MacBeath C E - - - 3M3xtaj
Table 1. The loci of $X_{k},k=1,7,8$ for 10 Poncelet conic configurations. ‘Inner’ and ‘Outer’ indicate if such conics are either ellipses (E) or circles (C). Under columns labeled $X_{k}$ ($k=1,7,8$), symbols ‘E’, ‘C’, ‘P’, ‘-’ indicate if the corresponding locus is an ellipse, a circle, a point, or a non-conic, respectively. The pattern that emerges is that $X_{7}$ and $X_{8}$ can only be conics if ‘Inner’ is a circle, in which case $X_{1}$ is a point. The confocal family is one exception: the inner conic is a (confocal) ellipse, and all three centers sweep conics. Note: the last column gives a hyperlink to an animated view of the family.
Note that for the case of the bicentric pair, the locus of both $X_{7}$ and $X_{8}$ are circles [21].
Figure 23. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are conics: (i) top left: confocal pair, live; (ii) bottom left: generic circular caustic, live; (iii) right: Chapple’s porism (bicentric $n=3$), live. Note that in (ii) and (iii), $X_{1}$ is stationary (center of caustic). Also note that in (iii), the loci of $X_{7}$ and $X_{8}$ are circles.
Figure 24. Three Poncelet families for which the loci of $X_{k}$, $k=1,7,8$ are non-conics: (i) top left: pair of homothetic ellipses, live; bottom left: pair of ‘dual’ ellipses, live; (iii) right: the MacBeath family, live.
7. Four special families
We study four special families of ellipse-inscribed Poncelet triangles about the incircle. Referring to Figure 25, these are: (i) focal-$X_{1}$: the caustic is centered on a focus of $E$; (ii) iso-$X_{2}$: caustic is on a first special point such that the barycenter $X_{2}$ is stationary; (iii) focal-$X_{4}$: caustic is centered on a second special point such that the orthocenter $X_{4}$ is stationary on a focus of $E$; (iv) iso-$X_{7}$: caustic is centered on a third special point such that the Gergonne point $X_{7}$ is stationary.
In the calculations below, $R$, $r$, $l_{i},i=1,2,3$ refer to a triangle’s circumradius, inradius, and sidelengths, respectively.
Figure 25. Four families of ellipse-inscribed Poncelet triangles about the incircle: focal-$X_{1}$ (top left); iso-$X_{2}$ (top right); focal-$X_{4}$ (bottom left); iso-$X_{7}$ (bottom right). Video
7.1. Focal-X(1)
Referring to Figure 25 (top left):
Proposition 23.
The caustic centered at $C_{1}=[c,0]$ closes Poncelet $N=3$ with radius $r_{1}$ given by:
$r_{1}={\frac{{b}^{2}}{{c}^{2}}\left(\sqrt{{a}^{2}+{c}^{2}}-a\right)}$
Proof.
Direct from Proposition 2.
Observation 9.
The focal-$X_{1}$ family is the polar image of the vertices of Chapple’s Porism (a porism of triangles interscribed between two circles [21]), with respect to the circumcircle, i.e., it is its tangential triangle. The former’s $X_{1}$ coincides with the latter’s circumcenter $X_{3}$.
While bicentric $n$-gons (a generalization of Chapple’s porism) conserve the sum of cosines [28], their polar family conserve $\sum_{i=1}^{n}{\sin(\theta_{i}/2)}$, for all $n\geq 3$ [2, prop.26].
Proposition 24.
For the focal-$X_{1}$ family, the invariant sum of half-angle sines is given by:
$\sum_{i=1}^{3}\sin\frac{\theta_{i}}{2}=\frac{c^{2}-a^{2}+a\sqrt{a^{2}+c^{2}}}{c^{2}}$
7.2. Iso-X(2)
Referring to Figure 25 (top right):
Proposition 25.
For a circular caustic with center and radius given by:
$C_{2}=\left[0,\frac{c\,b}{2a}\right],\,\,\,r_{2}=\frac{b}{2}$
The barycenter $X_{2}$ is stationary at
$\left[0,\frac{c\,b}{3a}\right]$
Corollary 11.
Over iso-$X_{2}$ triangles, the Nagel point $X_{8}$ is stationary at $E$’s center.
Proof.
Direct from the location of the stationary barycenter and the fact that $|X_{1}X_{8}|=3|X_{1}X_{2}|$ [37, Incenter, eqn.(11)]. In fact, $|X_{1}X_{8}|=3(c\,b/(2a)-c\,b/(3a))=c\,b/(2a)=|C_{2}|$. ∎
Corollary 12.
The Spieker center $X_{10}$, incenter of the medial triangle and the midpoint of $X_{1}$ and $X_{8}$ [18] will be stationary on $E$’s minor axis.
Observation 10.
Since $X_{2}$ is stationary, the family conserves $|X_{1}X_{2}|$, which for any triangle with sidelengths $l_{i}$ is given by [37, Triangle Centroid, eqn.(11)]:
$|X_{1}-X_{2}|^{2}=-\frac{\sum{l_{i}^{3}}+9\prod{l_{i}}-2\left(l_{2}l_{1}^{2}+l_{3}l_{1}^{2}+l_{2}^{2}l_{1}+l_{3}^{2}l_{1}+l_{2}l_{3}^{2}+l_{2}^{2}l_{3}\right)}{9\sum{l_{i}}}$
Lemma 3.
The barycenter $X_{2}$ of the MacBeath is stationary.
Proof.
The foci of the caustic are $X_{3}$ and $X_{4}$, therefore $X_{2}$ must be fixed, $|X_{3}X_{2}|=|X_{3}X_{4}|/3$ [37, Triangle Centroid, eqn. (13)]. Consider also that MacBeath family are the intouch triangles of bicentric polygons (Chapple’s porism). The $X_{2}$ of the intouch triangle is the Weill point $X_{354}$ of the reference [18], shown to be stationary over the bicentric family [21, Thm. 4.1]. ∎
Proposition 26.
Given an outer ellipse $E$ centered on $O$ and a point $C$ in its interior, there is an (elliptic) caustic $K$ centered on $C$ such that over Poncelet, $X_{2}$ on $OC$ is stationary.
Referring to Figure 26:
Proof.
Let $\mathcal{A}$ be an affine transform that sends $E$ to a circle $E^{\prime}$. Let $K^{\prime}$ be the unique MacBeath caustic in $E^{\prime}$ with a first focus at its center and the other the reflection of $E^{\prime}$ about $C^{\prime}=\mathcal{A}{\cdot}C$. Recall that conic centers are equivariant with respect to affinities. Let $K$ be the image of $K^{\prime}$ under $\mathcal{A}^{-1}$. The $X_{2}$ of the family in $(E,K)$ will be stationary because (i) the MacBeath’s centroid is stationary, Lemma 3 and (ii) the centroid is the only triangle center which is equivariant with respect to affinities [18]. $X_{2}$ must lie on $OC$ because affinities preserve collinearities. Notice it will nevertheless not lie on the major axis of $K$ (foci are not equivariant). ∎
Figure 26. left: To find a caustic at a given center $O_{c}$ that keeps $X_{2}$ stationary, find an affinity that sends the outer ellipse $E^{\prime}$to a circle $E^{\prime}$; right: in such an affine projection find the MacBeath caustic with a focus at the center and another at the reflection of the latter about the projection of $O_{c}$ under the affinity. The desired caustic on the left system is the image of said MacBeath caustic under the inverse affinity. Because affinities preserve collinearities, the stationary $X_{2}$ will like on $OO_{c}$, Video
7.3. Focal-X(4)
Referring to Figure 25 (bottom left):
Proposition 27.
The orthocenter $X_{4}$ is stationary at a focus $f=[\pm c,0]$ of $E$ for the circular caustic with center and radius given by:
$C_{4}=\left[\pm\frac{a^{2}c}{2a^{2}-c^{2}},0\right],\,\,\,r_{4}=\frac{a(a^{2}-c^{2})}{2a^{2}-c^{2}}$
Numeric search over the first forty-thousand triangle centers on [18] reveals:
Referring to Figure 27:
Observation 11.
For the focal-$X_{4}$ family the center (resp. other focus) of $E$ is $X_{7952}$ (resp. $X_{18283}$) on [18]. The locus of the circumcenter $X_{3}$ is an ellipse with a focus at $E$’s center. The locus of $X_{20}$ is twice as large, with a focus on $E$’s distal focus.
Figure 27. The focal-$X_{4}$ family: the outer ellipse $E$ has foci on $X_{4}$ and $X_{18283}$ and center on $X_{7952}$. Also shown are the loci of the orthocenter $X_{3}$, with a focus on said center, and that of the de Longchamps point $X_{20}$ (a reflection of $X_{4}$ on $X_{3}$), so twice as big, and with a focus on the other focus of $E$. Video
Definition 1.
A triangle’s polar circle is centered on $X_{4}$ and has squared radius given by [37, Polar Circle]:
$r_{pol}^{2}=4R^{2}-\frac{\sum{l_{i}^{2}}}{2}$
This quantity is positive (resp. negative) for obtuse (resp. acute triangles).
Proposition 28.
The focal-$X_{4}$ family conserves the (negative) squared radius of its polar circle at $r^{2}_{pol}=-b^{4}/(a^{2}+b^{2}).$
Proof.
The incenter-orthocenter squared distance is given by [37, Incenter, eqn.(6]:
(3) $|X_{1}-X_{4}|^{2}=2r^{2}+4R^{2}-\frac{\sum{l_{i}^{2}}}{2}=2r^{2}+r_{pol}^{2}$
Since for this family $|X_{1}X_{4}|$ is fixed as is the inradius $r$. The actual expression was obtained by CAS-based simplification. ∎
Referring to Figure 28 (left):
Definition 2.
The MacBeath Poncelet family is inscribed in a circle and circumscribes the MacBeath inconic, with foci at the circumcenter and its isogonal conjugate, the orthocenter. Its center is the Euler center $X_{5}$, the midpoint of $X_{3}$ and $X_{4}$ [37, MacBeath inconic].
For the purposes of this section, let $a,b$ denote the semiaxis’ lengths of the MacBeath inconic.
Lemma 4.
Over the MacBeath family, the sum of squared sidelengths, the sum of double-angle cosines, and the product of cosines are conserved. There are given by:
$\sum{l_{i}^{2}}$ $=9R^{2}-|X_{3}-X_{4}|^{2}=32a^{2}+4b^{2}$
$\sum{\cos(2\theta_{i})}$ $=\frac{c^{2}-3a^{2}}{2a^{2}}$
$\prod{\cos\theta_{i}}$ $=\frac{\sum{l_{i}^{2}}}{8R^{2}}-1=\frac{b^{2}}{8a^{2}}$
Proof.
The first relation is direct from [37, Circumcenter, eqn.(6)], noting that the (i) the circumradius $R$ is fixed and (ii) the caustic foci $X_{3},X_{4}$ are stationary. The second one is obtained via CAS-simplification. The third one is direct from [37, Circumradius, eqn.(5)]. ∎
Referring to Figure 28 (right):
Definition 3.
The dual Poncelet family is interscribed between two dual ellipses [36]. They are also homothetic if one is rotated by 90-degrees.
Figure 28. left: the ‘MacBeath’ family, live; right: the ‘dual’ family, live.
Proposition 29.
The (i) MacBeath and (ii) dual families conserve $r_{pol}^{2}<0$. These are given by:
$r_{pol,macbeath}^{2}$ $=\frac{|X_{3}-X_{4}|^{2}-R^{2}}{2}=-2b^{2}$
$r_{pol,dual}^{2}$ $=-\frac{a^{2}b^{2}}{a^{2}+b^{2}}$
Proof.
The first is direct from $|X_{3}-X_{4}|^{2}=9R^{2}-\sum{l_{i}^{2}}=2r_{pol}^{2}+R^{2}$ [37, Orthocenter, eqn.(14)], noting that the MacBeath circumradius $R$ is fixed and the caustic foci $X_{3}$ and $X_{4}$ are stationary. For the dual family, $X_{4}$ is stationary at the common center. Given the first relation it would therefore be sufficient to prove that $|X_{3}(t)|^{2}-R^{2}(t)$ is constant. The final expression is obtained via CAS simplification. ∎
Observation 12.
For the dual family the locus of the circumcenter $X_{3}$ is an ellipse homothetic to $E$ with factor $c^{2}/(2(a^{2}+b^{2}))$.
Conjecture 4.
A Poncelet triangle family maintains $X_{4}$ stationary only if the conics are configured as focal-$X_{4}$, MacBeath or Dual.
7.4. Iso-X(7)
Referring to Figure 25 (bottom right):
Proposition 30.
For the circular caustic with center and radius given by:
$C_{7}=\left[\frac{k_{7}}{2a},0\right],\,\,\,r_{7}=\frac{b^{2}}{2a}$
The Gergonne point will be stationary at:
$X_{7}=\left[\frac{2ak_{7}}{4a^{2}-b^{2}},0\right]$
where $k_{7}=\sqrt{4a^{4}-5a^{2}b^{2}+b^{4}}$.
Observation 13.
The iso-$X_{7}$ family is the polar image of the Brocard porism [25, 31] with respect to the circumcircle, i.e., its tangential triangle. The former’s stationary Gergonne point $X_{7}$ coincides with the stationary symmedian point $X_{6}$ of the Brocard porism [9].
Proposition 31.
The family conserves the sum of half-angle tangents, given by:
$\sum_{i=1}^{3}{\tan\frac{\theta_{i}}{2}}=\frac{\sqrt{4\,{a}^{2}-{b}^{2}}}{a}$
The $N>3$ generalization of the Brocard porism is known as the Harmonic family, studied [12].
Conjecture 5.
The sum of half-angle tangents is also conserved for the polar image of the Harmonic family, for all $N>3$.
Since $X_{7}$ is stationary, the quantity $|X_{1}X_{7}|$ will also be conserved. For a generic triangle with semiperimeter $s$, this is given by [23, Corollary 4.2]:
$|X_{1}-X_{7}|^{2}=r^{2}\left(1-\frac{3s^{2}}{(r+4R)^{2}}\right)$
Proposition 32.
Over Iso-$X_{7}$ triangles, this reduces to:
$|X_{1}-X_{7}|^{2}=\frac{b^{4}c^{2}}{4a^{2}(4a^{2}-b^{2})}$
Definition 4.
The Adams circle of a triangle is centered on $X_{1}$ and has radius $R_{A}$ given by [37, Adams Circle]:
$R_{A}=\frac{r\sqrt{\rho^{2}-l_{1}l_{2}l_{3}s-\rho s^{2}}}{\rho-s^{2}}$
where $\rho=l_{1}l_{2}+l_{2}l_{3}+l_{3}l_{1}$ and $s$ is the semi-perimeter.
Proposition 33.
The Iso-$X_{7}$ family conserves $R_{A}$ and its value is given by:
$R_{A}=\frac{b^{2}}{2a}\sqrt{\frac{5a^{2}-b^{2}}{4a^{2}-b^{2}}}$
Acknowledgements
We would like to thank Arseniy Akopyan, Clark Kimberling, Peter Moses, and Richard Schwartz, for commenting on our developing results, encouraging us constantly, and providing insights and formulas. The second author is fellow of CNPq and coordinator of Project PRONEX/ CNPq/ FAPEG 2017 10 26 7000 508.
Appendix A Affine triples
Triples $\alpha,\beta,\gamma$ used to express a given triangle center $X_{k}$ as the linear combination $\alpha X_{1}+\beta X_{2}+\gamma X_{3}$, where $\rho=r/R$, reproduced from [17, Table 1]:
$X_{1}$ $X_{2}$ $X_{3}$ $X_{4}$ $X_{5}$ $X_{7}$ $X_{8}$ $X_{9}$ $X_{10}$ $X_{11}$ $X_{12}$ $X_{20}$ $X_{21}$ $X_{35}$ $X_{36}$ $X_{40}$ $X_{46}$
$\alpha$ $1$ 00 00 00 00 $\frac{2\rho+4}{\rho+4}$ $-2$ $\frac{-\rho-2}{\rho+4}$ $-\frac{1}{2}$ $\frac{1}{1-2\rho}$ $\frac{1}{1+2\rho}$ 00 00 $\frac{1}{2\rho+1}$ $\frac{1}{1-2\rho}$ $-1$ $\frac{1+\rho}{1-\rho}$
$\beta$ 00 $1$ 00 $3$ $\frac{3}{2}$ $\frac{3\rho}{\rho+4}$ $3$ $\frac{6}{\rho+4}$ $\frac{3}{2}$ $\frac{-3\rho}{1-2\rho}$ $\frac{3\rho}{1+2\rho}$ $-3$ $\frac{3}{2\rho+3}$ 00 00 00 00
$\gamma$ 00 00 $1$ $-2$ $-\frac{1}{2}$ $\frac{-4\rho}{\rho+4}$ 00 $\frac{2\rho}{\rho+4}$ 00 $\frac{\rho}{1-2\rho}$ $\frac{-\rho}{1+2\rho}$ $4$ $\frac{2\rho}{2\rho+3}$ $\frac{2\rho}{2\rho+1}$ $\frac{-2\rho}{1-2\rho}$ $2$ $\frac{-2\rho}{1-\rho}$
$X_{55}$ $X_{56}$ $X_{57}$ $X_{63}$ $X_{65}$ $X_{72}$ $X_{78}$ $X_{79}$ $X_{80}$ $X_{84}$ $X_{90}$ $X_{100}$ $X_{104}$ $X_{119}$
$\alpha$ $\frac{1}{1+\rho}$ $\frac{1}{1-\rho}$ $\frac{2+\rho}{2-\rho}$ $\frac{-\rho-2}{\rho+1}$ $\rho+1$ $-\rho-2$ $\frac{\rho+2}{\rho-1}$ 1 $\frac{2\rho+1}{1-2\rho}$ $\frac{-\rho-2}{\rho}$ $\frac{-(\rho+1)^{2}}{\rho^{2}+2\rho-1}$ $\frac{2}{2\rho-1}$ $\frac{-2}{2\rho-1}$ $\frac{1}{2\rho-1}$
$\beta$ 00 00 00 $\frac{3}{\rho+1}$ 00 $3$ $\frac{-3}{\rho-1}$ $\frac{6\rho}{2\rho+3}$ $\frac{-6\rho}{1-2\rho}$ $\frac{6}{\rho}$ $\frac{6\rho}{\rho^{2}+2\rho-1}$ $\frac{-3}{2\rho-1}$ $\frac{3}{2\rho-1}$ $\frac{3\rho-3}{2\rho-1}$
$\gamma$ $\frac{\rho}{1+\rho}$ $\frac{-\rho}{1-\rho}$ $\frac{-2\rho}{2-\rho}$ $\frac{2\rho}{\rho+1}$ $-\rho$ $\rho$ 00 $\frac{-6\rho}{2\rho+3}$ $\frac{2\rho}{1-2\rho}$ $\frac{2\rho-4}{\rho}$ $\frac{2\rho(\rho-1)}{\rho^{2}+2\rho-1}$ $\frac{2\rho}{2\rho-1}$ $\frac{2\rho-2}{2\rho-1}$ $\frac{-\rho+1}{2\rho-1}$
$X_{140}$ $X_{142}$ $X_{144}$ $X_{145}$ $X_{149}$ $X_{153}$ $X_{165}$ $X_{191}$ $X_{200}$
$\alpha$ 00 $\frac{\rho+2}{2\rho+8}$ $\frac{-4\rho-8}{\rho+4}$ $\frac{4}{7}$ $\frac{-4}{6\rho-3}$ $\frac{4}{6\rho-3}$ $-\frac{1}{3}$ $-1$ $\frac{\rho+4}{\rho-2}$
$\beta$ $\frac{3}{4}$ $\frac{3\rho+6}{2\rho+8}$ $\frac{12-3\rho}{\rho+4}$ $\frac{3}{7}$ $\frac{-6\rho+9}{6\rho-3}$ $\frac{-6\rho-3}{6\rho-3}$ 00 $\frac{6}{2\rho+3}$ $\frac{-6}{\rho-2}$
$\gamma$ $\frac{1}{4}$ $\frac{-2\rho}{2\rho+8}$ $\frac{8\rho}{\rho+4}$ 00 $\frac{12\rho-8}{6\rho-3}$ $\frac{12\rho-4}{6\rho-3}$ $\frac{4}{3}$ $\frac{4\rho}{2\rho+3}$ 00
Appendix B Locus behaviors
Proposition 34.
Over Poncelet triangles with a circular caustic, from $X_{1}$ to $X_{1000}$, $X_{k}$ sweep a conic locus for the following $k$: 2, 3, 4, 5, 7, 8, 10, 11, 12, 20, 35, 36, 40, 42, 46, 55, 56, 57, 65, 79, 80, 81, 140, 145, 165, 171, 174, 176, 202, 213, 226, 354, 355, 365, 366, 367, 368, 370, 376, 381, 382, 388, 390, 396, 482, 484, 495, 496, 497, 498, 499, 506, 507, 546, 547, 548, 549, 550, 551, 553, 554, 559, 588, 590, 593, 597, 605, 609, 611, 612, 631, 632, 938, 939, 940, 941, 942, 944, 946, 950, 954, 955, 962, 975, 977, 980, 981, 989, 999, 1000.
Note that boldfaced entries above are triangle centers which are fixed combinations of $X_{2}$ and $X_{3}$ and will, over any Poncelet pair, sweep a conic locus [17].
Amongst some of the above-listed centers, six types of “locus behaviors” are observed:
  1. (1)
    $\mathcal{E}$-homothety: $X_{k}$, $k=2$ (see [30]), $4$, $8$ (concentric), $10$, $145$, $368$, $370$, $551$, $944$, $946$ are homothetic to $\mathcal{E}$.
  2. (2)
    Axis alignment: $X_{k}$, $k=7$, $79$, $390$, $1000$ are axis-aligned with $E$ but not homothetic to it.
  3. (3)
    C-focus: $X_{k}$, $k=5$, $12$, $355$, $495$, $496$ have $C$ for a focus.
  4. (4)
    C-major: the major axis of $X_{k}$, $k=3$ (foci on $E$’s axes), $35$, $40$, $55$, $165$ pass through $C$.
  5. (5)
    C-minor: the minor axis of $X_{k}$ $k=46$, $56$, $57$, $999$ pass through $C$.
  6. (6)
    Circle: $X_{k}$, $k=11$ (on incircle), $36$, $65$, $80$ (centered on C), $354$, $484$, $942$ are circles.
Pictures of the above can be found in [24].
Appendix C Fixed circumcircle family with equilateral
Referring to Figure 29, the family of contact triangles of a fixed incircle family is also Ponceletian: it is inscribed in a fixed circle and envelops the polar image of tangents to $\mathcal{E}$ with respect to the incircle, also a conic [1]. In particular, it is easy to see that the original family contains an equilateral triangle if and only if the family of contact triangles contains an equilateral. Thus we shall now study Poncelet families of unit circle-inscribed triangles and the many phenomena that arise when they contain an equilateral triangle.
Figure 29. The contact family $A^{\prime}B^{\prime}C^{\prime}$ is circle-inscribed and envelops a conic which is the polar image of tangents $\tau$ to $\mathcal{E}$ with respect to the incircle. The point $\tau^{\prime}$ indicated the instantaneous polarity.
Let $\mathcal{E}_{c}$ be an inconic with foci $f,g\in\mathbb{C}$, such that there is a Poncelet family of triangles inscribed in the unit circle $\mathbb{T}$ and circumscribing $\mathcal{E}_{c}$. Referring to Figure 30:
Lemma 5.
A Poncelet family of triangles inscribed in the unit circle and circumscribing an ellipse with foci $f,g\in\mathbb{C}$ contains an equilateral triangle iff $1/f+1/g$ is on the unit circle $\mathbb{T}$.
Proof.
First, assume that $T_{o}$ is an equilateral triangle in this family. Let $\zeta\in\mathbb{T}$ be one of the vertices of $T_{o}$ such that the other vertices are $e^{\frac{2\pi}{3}i}\zeta$ and $e^{\frac{4\pi}{3}i}\zeta$. Using the symmetric parameterization described in Theorem 1, let $\lambda_{o}\in\mathbb{T}$ be the parameter corresponding to the triangle $T_{o}$. Plugging in the vertices in the parameterization, we get
$f+g+\lambda_{o}\overline{f}\overline{g}=\zeta+e^{\frac{2\pi}{3}i}\zeta+e^{\frac{4\pi}{3}i}\zeta=\zeta\left(1+e^{\frac{2\pi}{3}i}+e^{\frac{4\pi}{3}i}\right)=0$
$\implies\lambda_{o}=-\frac{(f+g)}{\overline{f}\overline{g}}\implies\left|-{\frac{(f+g)}{\overline{f}\overline{g}}}\right|=|\lambda_{o}|=1$
$\implies|f+g|=|fg|\implies\left|\frac{1}{f}+\frac{1}{g}\right|=\left|\frac{f+g}{fg}\right|=1$
so $1/f+1/g$ is on the unit circle $\mathbb{T}$, as desired.
Conversely, suppose that $1/f+1/g$ is on the unit circle, that is, $|1/f+1/g|=1$. Rearranging, we have $|f+g|=|fg|$. Define $\lambda_{o}:=-(f+g)/(\overline{f}\overline{g})$. Then
$|\lambda_{o}|=\left|-\frac{(f+g)}{\overline{f}\overline{g}}\right|=\frac{|f+g|}{|\overline{f}\overline{g}}=\frac{|f+g|}{|fg|}=1$
By definition, $f+g+\lambda_{o}\overline{f}\overline{g}=0$. Moreover,
$fg+\lambda_{o}(\overline{f}+\overline{g})=\lambda_{o}\overline{\lambda_{o}}fg+\lambda_{o}(\overline{f}+\overline{g})=\lambda_{o}(\overline{\lambda_{o}}fg+\overline{f}+\overline{g})=$
$\lambda_{o}\overline{\left(f+g+\lambda_{o}\overline{f}\overline{g}\right)}=\lambda_{o}\overline{0}=0$
Hence, letting $z_{1},z_{2},z_{3}$ take the values of the 3 roots of $\zeta^{3}=\lambda_{o}$, we see that these $3$ roots form an equilateral triangle that satisfies the equations of the symmetric parameterization of Theorem 1. Thus, this equilateral triangle is part of the Poncelet family, as desired.
Figure 30. $T_{o}^{\prime}=A^{\prime}B^{\prime}C^{\prime}$ (brown) is an equilateral for which an inconic (green), centered at $\mathcal{O}$, is chosen, with foci at (complex) $f,g$. $ABC$ is a triangle (blue) in the Poncelet family defined by the circumcircle (black) and the chosen inconic. Over the family, $X_{110}$. the focus of the Kiepert parabola (not shown) is stationary at $(1/f+1/g)^{-1}$, as is its antipode $X_{74}$. Video
Henceforth, let $\mathcal{T}_{o}^{*}$ denote a Poncelet family of triangles interscribed between the unit circle and some inconic $\mathcal{E}_{c}$ satisfying Lemma 5.
Definition 5.
The Kiepert parabola of a triangle is tangent to the three sides of a triangle. Its focus is $X_{110}$, always on the circumcircle, and its directrix is the Euler line $X_{2}X_{3}$ [37].
Still referring to Figure 30, it follows from Lemma 5:
Corollary 13.
Over $\mathcal{T}_{o}^{*}$, $X_{110}$ is stationary at $(1/f+1/g)^{-1}$.
On [18], (i) the midpoint of $X_{3}$ and $X_{110}$ is called $X_{1511}$, and (ii) $X_{3233}$ is the vertex of the Kiepert inparabola [20]. Referring to Figure 31:
Observation 14.
Over $\mathcal{T}_{o}^{*}$, $X_{3233}$ sweeps a circle with diameter $X_{110}X_{1511}$, i.e., $1/2$.
Figure 31. A $T_{o}=ABC$ in $\mathcal{T}_{o}$ (blue) is shown (blue) along with its Kiepert parabola (orange), also an inconic, with focus at $X_{110}$. Over $\mathcal{T}_{o}$, the locus of its vertex $X_{3233}$ is a circle (dashed orange) whose diameter is $X_{110}X_{1511}$. Video
Referring to Figure 32, let $A_{eq}$ be one of the vertices of the equilateral in $\mathcal{T}_{o}^{*}$. Let $K$ be a chosen point on the circumcircle. Let $\mathcal{B}$ represent the external bisector of $\angle KX_{3}A_{eq}$.
It can be shown:
Proposition 35.
Over $\mathcal{T}_{o}^{*}$, $X_{110}$ will be stationary at $K$ for any $\mathcal{O}$ on the reflection of $\mathcal{B}$ about $X_{3}A_{eq}$.
Figure 32. The construction in Proposition 35: $K$ is chosen on the circumcircle; $A_{eq}$ is a vertex of the equilateral; $T_{o}=ABC$ is a Poncelet triangle in $\mathcal{T}_{o}^{*}$; $\mathcal{B}$ (dashed green) is the external bisector of $\angle KX_{3}A_{eq}$; $\mathcal{B}^{\prime}$ (dashed red) is its reflection about $X_{3}A_{eq}$; over all caustic centers $\mathcal{O}$ on $\mathcal{B}^{\prime}$, $X_{110}$ is stationary, over $\mathcal{T}_{o}^{*}$, at the chosen $K$.
Referring to Figure 33, let $T_{o}=ABC$ be some generic triangle. Let $\mathcal{O}$ be the center of an inconic $\mathcal{E}_{c}$. Let $\mathcal{T}_{o}$ be the Poncelet family defined by the circumcircle and $\mathcal{E}_{c}$. Let $\mathcal{L}_{35}$ be the perpendicular bisector of $X_{3}X_{5}$.
It can be shown:
Proposition 36.
$\mathcal{T}_{o}$ will contain an equilateral iff $\mathcal{O}$ lies on $\mathcal{L}_{35}$. Furthermore, the stationary $X_{110}$ is independent of $\mathcal{O}$.
Figure 33. The construction of Propositions 36 and 37. $X_{110}$ (of $T_{o}=ABC$) is stationary over Poncelet if $\mathcal{O}$ lies on $\mathcal{L}_{35}$ (dashed red), the perpendicular bisector of $X_{3}$ and $X_{5}$. A vertex $A_{eq}$ of the equilateral contained in the family is shown.
Let $\alpha=\angle X_{110}X_{3}\mathcal{O}$. Let $\mathcal{O}^{\prime}$ be the counterclockwise rotation of $\mathcal{O}$ about $X_{3}$ by $\alpha/3+\pi$. Still referring to Figure 33.
It can be shown:
Proposition 37.
One vertex of the equilateral contained in $\mathcal{T}_{o}$ is located at the intersection of $X_{3}\mathcal{O}^{\prime}$ with the circumcircle.
Let $\mathcal{T}_{o}^{\prime}$ be the family of tangential triangles to $\mathcal{T}_{o}$, i.e., whose sides are tangent to the circumcircle at the vertices of $\mathcal{T}_{o}$ [37, Tangential Triangle]. Clearly, this family circumscribes a circle and will be inscribed in an ellipse $\mathcal{E}$, called above $\mathcal{T}$.
Observation 15.
$\mathcal{E}$ is the polar image of the chosen inconic with respect to the circumcircle.
Corollary 14.
Since the tangential of an equilateral is an equilateral, $\mathcal{T}_{o}^{\prime}$ will contain an equilateral, and therefore the center of its incircle will be on $\mathcal{E}_{eq}$.
In turn, this leads to a restatement of Proposition 15:
Corollary 15.
The Feuerbach point $X_{11}$ of $\mathcal{T}_{o}^{\prime}$ will be stationary.
Proof.
$X_{110}$ of a reference triangle is $X_{11}$ of its tangential. Conversely, $X_{11}$ of the reference is $X_{110}$ of the contact triangle. ∎
On [18], $X_{65}$ is the orthocenter $X_{4}$ of the contact triangle.
Corollary 16.
The locus of $X_{65}$ of the tangential family is a circle.
Proof.
The contact family is Poncelet (sides envelop the polar image of $\mathcal{E}$ wrt incircle) and circle-inscribed. $X_{4}$ is always homothetic to a 90-degree-rotated version of the inscribing conic. ∎
Appendix D Semiaxes of the X(7) locus
Let $(xc,yc)$ represent the center of the circular caustic. The square of the major semiaxis length $a_{7}$ of the $X_{7}$ locus is given by:
(a*b^2*(16*a^17*(b^4+22*b^2*yc^2+9*yc^4)+16*a*b^12*xc^2*(b^2-yc^2)*(28*b^4-30*xc^2*yc^2+28*yc^4-b^2*(51*xc^2+56*yc^2))+8*a^15*(35*b^6-36*xc^2*yc^4-2*b^4*(4*xc^2+175*yc^2)-3*b^2*(44*xc^2*yc^2+119*yc^4))+a^13*(-2559*b^8+144*xc^4*yc^4+b^6*(-812*xc^2+2517*yc^2)+3*b^4*(32*xc^4+2168*xc^2*yc^2+4945*yc^4)+3*b^2*(352*xc^4*yc^2+812*xc^2*yc^4+189*yc^6))+a^3*b^10*(64*b^8+60*xc^6*yc^2-1212*xc^4*yc^4+832*xc^2*yc^6+64*yc^8-64*b^6*(73*xc^2+4*yc^2)+12*b^4*(573*xc^4+848*xc^2*yc^2+32*yc^4)-b^2*(141*xc^6+6960*xc^4*yc^2+6336*xc^2*yc^4+256*yc^6))+a^11*b^2*(2698*b^8+b^6*(5397*xc^2+18490*yc^2)+18*b^4*(42*xc^4-983*xc^2*yc^2-1409*yc^4)-4*xc^2*yc^2*(88*xc^4-105*xc^2*yc^2+81*yc^4)-b^2*(64*xc^6+4608*xc^4*yc^2+15699*xc^2*yc^4+2226*yc^6))+a^5*b^8*(112*b^8+8*b^6*(1545*xc^2-22*yc^2)-3*b^4*(4991*xc^4+11960*xc^2*yc^2+48*yc^4)+b^2*(138*xc^6+22881*xc^4*yc^2+23064*xc^2*yc^4+368*yc^6)+4*(3*xc^8+18*xc^6*yc^2-75*xc^4*yc^4+114*xc^2*yc^6-40*yc^8))-a^7*b^6*(3876*b^8+b^6*(383*xc^2-6744*yc^2)-6*b^4*(1897*xc^4+6571*xc^2*yc^2-288*yc^4)+3*b^2*(-85*xc^6+9530*xc^4*yc^2+14813*xc^2*yc^4+424*yc^6)+4*(7*xc^8+171*xc^6*yc^2-756*xc^4*yc^4+483*xc^2*yc^6-33*yc^8))+a^9*b^4*(9097*b^8-b^6*(14218*xc^2+26815*yc^2)+b^4*(-3105*xc^4+300*xc^2*yc^2+14943*yc^4)+b^2*(-196*xc^6+15861*xc^4*yc^2+39894*xc^2*yc^4+2811*yc^6)+4*(4*xc^8+226*xc^6*yc^2-639*xc^4*yc^4+354*xc^2*yc^6-9*yc^8))-192*b^11*xc^2*(b^2-yc^2)^2*d-128*a^14*b*(b^2+3*yc^2)*d+192*a^12*b*(2*b^4+4*xc^2*yc^2+3*yc^4+b^2*(2*xc^2+23*yc^2))*d-4*a^2*b^9*(80*b^6-75*xc^4*yc^2+144*xc^2*yc^4-80*yc^6-120*b^4*(7*xc^2+2*yc^2)+3*b^2*(79*xc^4+232*xc^2*yc^2+80*yc^4))*d+12*a^10*b*(258*b^6-32*xc^4*yc^2+33*xc^2*yc^4-b^4*(79*xc^2+1116*yc^2)-2*b^2*(16*xc^4+193*xc^2*yc^2+211*yc^4))*d+12*a^4*b^7*(168*b^6+5*xc^6-99*xc^4*yc^2+178*xc^2*yc^4-48*yc^6-2*b^4*(577*xc^2+192*yc^2)+b^2*(263*xc^4+1320*xc^2*yc^2+264*yc^4))*d-12*a^6*b^5*(47*b^6+15*xc^6-90*xc^4*yc^2+102*xc^2*yc^4-25*yc^6-b^4*(1130*xc^2+799*yc^2)+b^2*(232*xc^4+2404*xc^2*yc^2+777*yc^4))*d-4*a^8*b^3*(2579*b^6-32*xc^6-54*xc^4*yc^2+129*xc^2*yc^4+9*yc^6+3*b^4*(31*xc^2-447*yc^2)-3*b^2*(62*xc^4+1570*xc^2*yc^2+949*yc^4))*d))/((a^2-b^2)^2*(16*a^6-9*b^4*xc^2-8*a^4*(b^2+2*xc^2)+a^2*b^2*(b^2+24*xc^2-yc^2))*(-(a^2*b^2*(8*b^2+xc^2-24*yc^2))+a^4*(b^2-9*yc^2)+16*(b^6-b^4*yc^2))^2)
where: d=sqrt((a^4-a^2*xc^2+b^2*xc^2)*(b^4+a^2*yc^2-b^2*yc^2)). The expression for the minor semi-axis $b_{7}$ is the above where $(a,b)$ and $(x_{c},y_{c})$ are permuted.
Appendix E Table of symbols
symbol meaning
$\mathcal{E},\mathcal{E}_{c}$ Poncelet conics (incidence and tangency)
$a,b,c$ semi-axis’ and half-focal lengths of $\mathcal{E}$
$a_{c},b_{c},c_{c}$ idem for $\mathcal{E}_{c}$
$\mathcal{K},C=(x_{c},y_{c})$ incircle and its center (= $X_{1}$)
$T,l_{i},\theta_{i}$ a Poncelet triangle, sidelengths, and internal angles
$r,R$ inradius and circumradius of a $T$
$\mathcal{E}_{eq},a_{eq},b_{eq}$ (elliptic) locus of $\mathcal{E}$-inscr. equilat. centroids and semi-axes
$X_{k}$ triangle center indexed in Kimberling’s [18]
$X_{1}$ Incenter: meet of angle bisectors
$X_{2}$ Barycenter: meet of medians
$X_{3}$ Circumcenter: meet of perpendicular bisectors
$X_{4}$ Orthocenter: meet of altitudes
$X_{5}$ Center of Euler’s (9-pt) circle
$X_{7}$ Gergonne point: perspector of the contact triangle
$X_{8}$ Nagel’s point: perspector of the extouch triangle
$X_{11}$ Feuerbach’s point: touchpoint of incircle and Euler’s circle
$X_{36}$ circumcircle-inverse of $X_{1}$
$X_{59}$ isogonal conjugate of $X_{11}$
$X_{80}$ reflection of $X_{1}$ on $X_{11}$
$C_{i}=(x_{i},y_{i}),a_{i},b_{i}$ center and semiaxes of the (elliptic) locus of $X_{i}$’s
$F_{i},F_{i}^{\prime}$ foci of the (elliptic) locus of $X_{i}$
$X_{i}^{*}$ stationary $X_{i}$
$\mathcal{T}$ Poncelet triangle family about the incircle
$\mathcal{L}_{k}$ locus of $X_{k}$ over $\mathcal{T}$
$\mathcal{T}^{*}$ the family $\mathcal{T}$ containing an equilateral
$\mathcal{T}_{g}$ Poncelet triangle interscribed between two generic conics
$\mathcal{T}_{o}$ circle-inscribed Poncelet triangle family
$\mathcal{T}_{o}^{*}$ the family $\mathcal{T}_{o}$ containing and equilateral
$\mathcal{T}_{o}^{\prime}$ tangential triangles to $\mathcal{T}_{o}^{*}$, identical to $\mathcal{T}^{*}$
$f,g$ foci of caustic of $\mathcal{T}_{o}$
$\lambda$ parameter of symmetric parametrization [4]
Table 2. Symbols used in the article. For geometric constructions for the $X_{i}$, see [37].